### v1

# Multi-scale Cascaded Large-Model for Whole-body ROI Segmentation

Abstract
Organs-at-risk segmentation is critical for ensuring the safety and precision of radiotherapy and surgical procedures. However, existing methods for organs-at-risk image segmentation often suffer from uncertainties and biases in target selection, as well as insufficient model validation experiments, limiting their generality and reliability in practical applications. To address these issues, we propose an innovative cascaded network architecture called the Multi-scale Cascaded Fusing Network (MCFNet), which effectively captures complex multi-scale and multi-resolution features. MCFNet includes a Sharp Extraction Backbone and a Flexible Connection Backbone, which respectively enhance feature extraction in the downsampling and skip-connection stages. This design not only improves segmentation accuracy but also ensures computational efficiency, enabling precise detail capture even in low-resolution images. We conduct experiments using the A6000 GPU on diverse datasets from 671 patients, including 36,131 image-mask pairs across 10 different datasets. MCFNet demonstrates strong robustness, performing consistently well across 10 datasets. Additionally, MCFNet exhibits excellent generalizability, maintaining high accuracy in different clinical scenarios. We also introduce an adaptive loss aggregation strategy to further optimize the model training process, improving both segmentation accuracy and efficiency. Through extensive validation, MCFNet demonstrates superior performance compared to existing methods, providing more reliable image-guided support. Our solution aims to significantly improve the precision and safety of radiotherapy and surgical procedures, advancing personalized treatment. The code has been made available on GitHub: https://github.com/Henry991115/MCFNet.
Index Terms:
Whole-body ROI segmentation, multi-scale cascaded Large-Model, linear attention transformer, adaptive loss aggregation strategy.
I Introduction
Medical image segmentation aims to quickly and accurately separate different tissues, organs, or lesion areas, facilitating clinical applications such as disease analysis, lesion detection, and treatment planning [1, 2]. In the field of medical image analysis, accurately identifying and segmenting malignant tumors is crucial for early cancer diagnosis, treatment planning, and efficacy evaluation. However, malignant tumors are often complex in shape, vary in size, have blurred boundaries, and are closely connected to surrounding tissues, posing significant challenges for accurate tumor identification and segmentation [3, 4, 5]. Accurately identifying and locating organs-at-risk (OAR) can help doctors and medical teams plan and execute treatment plans. In radiation therapy, precise identification of OAR can assist doctors in minimizing the impact of treatment on surrounding healthy tissues, thereby reducing adverse effects caused by the treatment [6, 7]. Tumor and OAR segmentation are fundamental to treatment planning in radiation therapy. Tumor segmentation helps doctors determine the target area for accurately aiming the radiation dose. Meanwhile, OAR segmentation allows doctors to identify and protect surrounding healthy tissues, avoiding unnecessary damage during treatment [8, 9].
With the emergence of deep learning models, traditional methods constrained by human, financial, and time costs are gradually being abandoned [10, 11, 12, 13]. Deep learning-based models can automatically learn features from images and efficiently extract features of complex structures and textures, enabling the identification of ROI contours and facilitating subsequent medical image analysis [14, 15]. Most current efficient segmentation models require a substantial amount of high-performance GPU computing resources for training and have a strong dependency on large-scale annotated data. However, obtaining high-quality annotated data requires a significant amount of time and human resources [16, 17].
Most current deep learning models for segmentation are based on UNet [18] and Transformer [19] architectures. UNet is renowned for its efficient and simple structure, which can effectively capture local features in images, while Transformers are known for their outstanding ability to capture global information [20, 21, 22]. The combination of UNet’s local feature extraction capabilities with the Transformer’s global information capture ability allows segmentation models to consider both local details and overall structure in images, thereby improving segmentation accuracy and robustness. This combination provides a more comprehensive and efficient solution for medical image analysis, offering strong support for advancements in the medical imaging field. However, the segmentation performance of most current methods is limited by the application scenario, and the robustness and generalization abilities of the models are usually weak. High-performance models like TransUNet [23], Swin-Unet [24], and MISSFormer [25] have demonstrated their capabilities primarily in segmenting eight organs in the heart and abdominal regions. UCTransNet [26] has expanded its tasks to include the segmentation of glands and cell nuclei. Nevertheless, in our actual experiments, we found that the segmentation performance of these methods varies significantly when applied to datasets with different image characteristics.
Existing works primarily focus on segmenting certain organs-at-risk and lesions, but they cannot segment multiple regions of the entire body, highlighting the limited application scenarios and the relatively poor robustness and generalization capabilities of these models. To address this, we propose conducting experiments on datasets that cover multiple organs-at-risk and tumors across the entire body. To enhance the model’s generalization ability for different image characteristics, we intend to capture image features from multiple angles using multi-scale and multi-resolution images, enabling the model to better adapt to current application scenarios. Finally, we design the FCB (Flexible Connection Backbone), which is targeted for multi-region segmentation, as one of the strong backbones. Additionally, we design a SEB (Sharp Extraction Backbone) to assist the FCB in better feature extraction. Cascading the two backbones can ensure high segmentation performance while designing the model to enhance its robustness and generalization ability. Our contributions are summarized as follows:
  • $\bullet$
    Novel Multi-scale Cascaded Large-Model: We propose a novel Multi-scale Cascaded U-shaped Network for whole-body ROI segmentation. This network cascades two backbone networks, namely our designed FCB (Flexible Connection Backbone) and SEB (Sharp Extraction Backbone). The two backbones are used to handle images of different scales and resolutions, with feature fusion occurring in the four skip connections and the down-sampling bottom layers of the network.
  • $\bullet$
    Innovative Adaptive Loss Aggregation Strategy: We introduce a novel adaptive loss aggregation strategy named Adaptive-MFA (Adaptive Multi-scale Feature-Mixing Loss Aggregation). By using this strategy, we can create new prediction maps by combining available prediction maps, thereby optimizing the model training as much as possible.
  • $\bullet$
    Extensive Validation for Robustness and Generalization: To thoroughly validate the model’s robustness and generalization capabilities, we selected ten datasets targeting organs-at-risk and tumors throughout the body for experiments. The segmented regions include the brain, thoracic cavity, abdomen, and lower limbs, with two of these datasets provided by our collaborating hospital.
II Related Work
In this section, we introduce commonly used architectures for medical image segmentation as well as cascade network architectures.
II-A Medical Image Segmentation
Medical image segmentation refers to accurately separating different structures (such as organs, lesion regions, etc.) in medical images, which is a crucial task in medical image analysis and computer-aided diagnosis. To address this task, research on U-shaped networks based on encoder-decoder structures has seen explosive growth. UNet innovatively introduced the encoder-decoder architecture, enabling the aggregation of features from different stages through skip connections. Subsequently, UNet++ [28] and UNet 3+ [29] improved the skip connection mechanism, further enhancing the effectiveness of feature aggregation.
In recent years, the application of Transformers in medical image segmentation has attracted widespread attention. Originally used in the field of natural language processing, Transformers have shown great potential in computer vision tasks due to their powerful modeling capabilities and flexibility. Compared to traditional convolutional neural networks (CNNs), Transformer architectures primarily rely on self-attention mechanisms, allowing them to capture long-range dependencies in images globally. ViT [30] is one of the first models to apply Transformers to image classification tasks, processing images by splitting them into fixed-size patches and treating these patches as sequences inputted to the Transformer for processing. The same concept has been introduced into medical image segmentation, such as the TransUNet model, which significantly improves segmentation performance by combining the local feature extraction capability of UNet with the global contextual modeling capability of Transformers. Swin Transformer is a hierarchical Transformer architecture that further enhances computational efficiency and feature extraction capability through a sliding window mechanism. Swin-Unet applies this architecture to medical image segmentation tasks and achieves excellent results. Recently, Rahman et al. proposed the Cascaded MERIT [31] model, which effectively improves the accuracy and stability of medical image segmentation through techniques such as multiscale hierarchical feature extraction and cascaded attention decoding.
II-B Cascaded Networks for Medical Image Segmentation
In recent years, cascade networks have gained considerable attention in medical image segmentation. These networks concatenate multiple modules, each handling specific subtasks, gradually improving overall segmentation performance. Cascade networks can integrate feature information at different scales and decompose tasks into multiple stages for optimization, with each stage utilizing feature representations from the preceding stage [31]. Information propagation is achieved through mechanisms such as skip connections and attention mechanisms. In experimental validation, cascade networks have demonstrated effectiveness and superiority in medical image segmentation tasks, offering significant potential as diagnostic and therapeutic tools in clinical medicine.
Recently, the CASCADE model introduced a novel attention-based decoder that learns local (contextual) relationships between pixels, addressing the shortcomings of Transformer models in learning local relationships. G-CASCADE [32] refines multi-stage feature maps generated by a hierarchical Transformer encoder through graph convolutional blocks, contributing to improved accuracy and robustness of segmentation results through progressive refinement. The Cascaded MERIT [31] model improves generalization capability by introducing a multiscale self-attention mechanism. Traditional single-scale self-attention mechanisms may not effectively capture multiscale information in images, whereas Cascaded MERIT computes self-attention at multiple scales to comprehensively understand image features, thereby enhancing model performance. MCANet proposes a multi-scale cross-axis attention mechanism that efficiently captures information from different sizes and shapes of lesion regions or organs in medical images. By computing dual cross attentions between two parallel axial attentions, better global information capture is achieved, facilitating accurate segmentation of target areas in medical images.
III Proposed Method
In this section, we introduce the multi-scale cascade network model MCFNet designed for whole-body ROI segmentation. First, we provide an overall introduction to the model, then we describe each component of the model in detail, and finally, we explain the adaptive loss aggregation strategy we designed.
Figure 1: Illustration of the overall architecture of MCFNet.
III-A MCFNet for Whole-body ROI Segmentation
Fig. 1 shows the overall architecture of our designed MCFNet network. MCFNet consists of three parts: Cascaded Backbones, Decoders with cascaded skip connections, and Aggregation. The Cascaded Backbones module is composed of SEB (Acute Extraction Backbone) and FCB (Elegant Connection Backbone), enabling multi-scale and multi-resolution input for the model. The Decoders with cascaded skip connections module mainly achieves feature aggregation at skip connections and bridge layers in the cascade network, enhancing feature representation by fully utilizing feature information from different levels and different sub-networks. The Aggregation module primarily reflects the aggregation of the decoder output layers and the application of our designed adaptive loss aggregation strategy.
III-B Cascaded Backbones in MCFNet
The cascaded backbone network consists of SEB (Sharp Extraction Backbone) and FCB (Flexible Connection Backbone). These two sub-backbone networks are responsible for feature extraction from images of different resolutions. Generally, high-resolution images can provide richer details and information, aiding the model in more accurately identifying and segmenting structures and regions within the images. However, this increases computational complexity and resource requirements. Low-resolution images are typically more robust to noise and interference but lose detailed information. Considering application requirements and computational resources, we choose to use FCB for processing $224\times 224$ images and SEB for assisting with the processing of $256\times 256$ images. In the encoder stage, the FCB sub-network uses traditional max pooling and convolution operations. This means that FCB has some limitations in feature extraction during the downsampling stage. Therefore, the design of SEB focuses on feature extraction during the downsampling stage while ensuring a relatively small number of parameters and computational load.
Specifically, SEB enhances the recognition and extraction of detailed features during the downsampling stage for high-resolution images by adding SE attention mechanisms to the encoding layers. As the feature extraction process deepens in SEB, the images go through BatchNorm, MaxPool, and convolution operations via the SEB Stem and encoding layers. The prediction map of SEB is finally output from the highest layer of the SEB decoder and is subjected to a Hadamard product operation with the input image of SEB to obtain a new output image. This new output image is then resized to the $224\times 224$ input image required by FCB. The low-resolution images are input into FCB and processed through the FCB Stem and encoding layers. The four layers of feature maps are output through a skip connection mechanism and subsequently aggregated with those from SEB.
III-C Cascaded Skip-connection Module in MCFNet
As shown in Fig. 2, this section mainly introduces the cascaded skip-connection mechanism. The efficient backbone model, FCB, is used to ensure the effectiveness of feature extraction, while a lightweight and simple SEB is employed to capture detailed information in high-resolution images, thus compensating for the limitations of FCB.
Figure 2: (a) Illustration of the CSM (Cascaded Skip-connection Module). (b) Illustration of the LAT (Linear Attention Transformer Block) in the CSM.
Specifically, when the encoding layers of FCB reach the lowest level after downsampling, they enter the CAB (Cross-Attention Bridge Layer). This module compensates for the loss of low-level features and then aggregates features with the output feature maps of the fourth encoding layer in SEB. For skip connections in FCB, the features first pass through the LAT (Linear Attention Transformer Block), enabling comprehensive and efficient feature extraction across both channel and spatial dimensions. As shown in Fig. 2, the structure of LAT is based on our published work, FSCA-Net, specifically the PAT (Parallel Attention Transformer Block). We make new improvements to accelerate the convergence speed of the transformer-like structure and reduce the number of model parameters caused by its architecture. After the input tensor enters the LAT module, we first use a linear layer for dimensionality reduction to align the tensor dimensions with the channel numbers. The compressed dimensional feature maps are then processed sequentially through layers including LN, 2D-SCCA attention, and another LN layer. Finally, we obtain the output features by passing the data through a linear layer for dimensional restoration.
Subsequently, the feature maps outputted by LAT are aggregated with the feature maps outputted by the skip connections of resized SEB. The aggregated feature maps are then fed into the decoding layer of FCB through skip connections, which directly propagate shallow features to deeper layers, preserving fine-grained information captured by shallow networks. In SEB, the output prediction map ($S_{output}$) from the highest decoding layer is fed into the SAM (Sigmoid Activation Map Module). This module mainly utilizes a 1x1 convolution to output a single-channel feature map, and then applies the $Sigmoid$ activation function to nonlinearly reweight features, enhancing the representation of key features while suppressing unimportant ones. The calculation formula for the SAM module is as follows:
$Y(i,j)=w\cdot X(i,j)+b,$ (1)
$Z(i,j)=\sigma(Y(i,j))=\frac{1}{1+{e}^{-Y(i,j)}},$ (2)
where the weight is denoted as $W$, the input image is $X$ (with dimensions $H\times W$), the output of the convolution operation is $Y$ (with dimensions $H\times W$), the bias term is denoted as $b$, the output of the Sigmoid activation function is $Z$ (with dimensions $H\times W$), and the values of the output image $Z$ are constrained between 0 and 1. After the $S_{output}$ undergoes the SAM operation, the final output image $Z$ is obtained. This output image is then subjected to Hadamard product operation with the input image $S_{input}$ of SEB, followed by a resize operation, as shown in the following formula:
$\hat{O}=S_{input}(i,j)\cdot Z(i,j),$ (3)
where $\hat{O}$ is the output image obtained from the Hadamard product. The resize operation is implemented using bilinear interpolation. The bilinear interpolation is defined as follows:
$\begin{split}f(x,y)=&(1-\alpha)\cdot(1-\beta)\cdot pixel(i_{1},j_{1})\\ &+\alpha\cdot(1-\beta)\cdot pixel(i_{2},j_{2})\\ &+(1-\alpha)\cdot\beta\cdot pixel(i_{3},j_{3})\\ &+\alpha\cdot\beta\cdot pixel(i_{4},j_{4}),\end{split}$ (4)
Figure 3: Illustration of the Aggregation module.
where the target image coordinates are denoted as $(x,y)$, the original image coordinates are $(i,j)$, and the four nearest neighbor pixels in the original image are $(i_{1},j_{1})$, $(i_{2},j_{2})$, $(i_{3},j_{3})$, $(i_{4},j_{4})$. Additionally, $\alpha=x-i_{1}$, $\beta=y-j_{1}$, and $pixel(i,j)$ represent the pixel value at coordinates $(i,j)$ in the original image. Each decoding layer of SEB and FCB generates a prediction map, representing the predicted results for the input. These prediction maps come from different scales and levels, reflecting different levels of feature information and semantic understanding. We feed these prediction maps into the Aggregation module for the aggregation of multi-scale information, aiming to comprehensively utilize information from different scales and improve the accuracy and robustness of the final prediction results.
III-D Aggregation in MCFNet
As shown in Fig. 3, the module mainly summarizes the process of multi-scale feature aggregation in the decoding layer stage of two sub-networks, as well as the application of our newly designed adaptive loss aggregation strategy. Multi-scale feature aggregation helps the model utilize both global contextual information and local details, assisting the model in capturing finer image details and complex edge structures. This is particularly important for handling complex small structures in medical images, such as blood vessels and tumor boundaries. Moreover, merging features of different scales helps generate richer and more diverse feature representations, enabling the model to capture more comprehensive image information, effectively dealing with noise and deformation in images, and enhancing the segmentation capability for objects of different sizes. This allows the model to perform more stably across different types of medical images.
Figure 4: Flowchart of the Adaptive-MFA Strategy.
Specifically, the predicted maps output from the four layers of the encoding stage of SEB and B are pairwise fed into the corresponding convolutional blocks for channel-wise processing, obtaining feature maps with the desired specified number of channels. The calculation formula for this part is shown below:
$O_{i}=Conv_{(C_{i},N,K)}(E_{i})(i=1,2,3,4),$ (5)
where $E$ refers to the input feature map, $O$ refers to the output feature map, $C$ is the number of input channels, $N$ is the number of label classes, and $K$ is the convolution kernel size. Additionally, $C_{1}=64$, $C_{2}=64$, $C_{3}=128$, $C_{4}=256$. After the convolution block operations, the two sub-networks use bilinear interpolation resizing to match the dimensions of the output feature maps from FCB and SEB, followed by feature map aggregation. Subsequently, the four layers of different scale feature maps are fed into our designed adaptive loss aggregation strategy for prediction map data augmentation, thereby optimizing the training of the entire model. Through the adaptive search optimization accompanying the entire training process, the four layers of feature maps are aggregated to obtain the final output prediction map $Pred$ of the MCFNet model. The aggregation of the four layers of feature maps can be expressed by the following mathematical formula:
$Pred=u\times p_{1}+v\times p_{2}+w\times p_{3}+x\times p_{4},$ (6)
where $p_{1}$, $p_{2}$, $p_{3}$, $p_{4}$ represent the feature maps from the four prediction heads, and $u$, $v$, $w$, $x$ are the corresponding weight coefficients for the four feature maps. In our experiments, we set all the weight coefficients to 1. In the following section, we will provide a detailed introduction to the adaptive loss aggregation strategy.
III-E Adaptive-MFA: Adaptive Multi-scale Feature-Mixing Loss Aggregation
To better optimize model training, we have designed an adaptive loss aggregation strategy. The main idea is to create new prediction maps by combining available prediction maps. By integrating various prediction maps, the model can better handle errors in different situations. Adaptively selecting and weighting different prediction maps can optimize the final prediction results based on performance in different scenarios. This allows the model to better capture complex patterns in the data, thereby improving prediction accuracy.
The entire strategy flowchart is illustrated in the Fig. 4. Specifically, we randomly combine the output prediction maps $P_{1}$, $P_{2}$, $P_{3}$ and $P_{4}$ from the four layers of the decoder stage, resulting in a total of 15 non-empty subset combinations. Subsequently, we divide the 15 combinations into four sets $S_{1}$, $S_{2}$, $S_{3}$ and $S_{4}$, and assign corresponding weights $W_{1}$, $W_{2}$, $W_{3}$ and $W_{4}$ to these sets. These weight coefficients are adaptively adjusted during model training. Regarding the setting of weight coefficients, since $W_{1}+W_{2}+W_{3}+W_{4}=1$ imposes a competitive relationship among the four weight coefficients, it restricts the optimization of model training as it cannot account for all prediction map combinations. In other words, some prediction maps might contribute minimally to the optimization of the training loss but are still crucial for the model training process. Therefore, we discard the $W_{1}+W_{2}+W_{3}+W_{4}=1$ restriction and only stipulate that the initial values satisfy $W_{1}=W_{2}=W_{3}=W_{4}$. The specific contents of the four sets are as follows:
$S_{1}=\Big{\{}\{p_{1}\},\{p_{2}\},\{p_{3}\},\{p_{4}\}\Big{\}},$ (7)
$\begin{split}S_{2}=&\Big{\{}\{p_{1},p_{2}\},\{p_{1},p_{3}\},\{p_{1},p_{4}\},\\ &\{p_{2},p_{3}\},\{p_{2},p_{4}\},\{p_{3},p_{4}\}\Big{\}},\end{split}$ (8)
$\begin{split}S_{3}=&\Big{\{}\{p_{1},p_{2},p_{3}\},\{p_{1},p_{2},p_{4}\},\\ &\{p_{1},p_{3},p_{4}\},\{p_{2},p_{3},p_{4}\}\Big{\}},\\ \end{split}$ (9)
$S_{4}=\Big{\{}\{p_{1},p_{2},p_{3},p_{4}\}\Big{\}}.$ (10)
We perform feature map aggregation operations according to the combinations of each subset within each set. Subsequently, we calculate the total loss for each of the four sets, denoted as $L_{1}$, $L_{2}$, $L_{3}$ and $L_{4}$. By incorporating the corresponding weight coefficients, we obtain the final total loss, $LOSS$, which is expressed as:
$LOSS=W_{1}*L_{1}+W_{2}*L_{2}+W_{3}*L_{3}+W_{4}*L_{4},$ (11)
the $LOSS$ obtained after each epoch during the training phase is fed back to the model itself, optimizing the model training and adaptively adjusting the weights. This process continues throughout the entire model training phase.
In summary, randomly combining and aggregating the outputs of the four-layer decoder allows full utilization of multi-level features, enhancing the model’s predictive capability, robustness, and flexibility. This approach also provides new ideas and methods for model design and optimization. Furthermore, through experiments, we find that the weight coefficients of prediction maps tend to favor different sets across different datasets. Therefore, adaptively searching for the optimal weight coefficients better meets the demands of different scenarios, thereby improving the model’s generalization ability and prediction accuracy. This adaptive search method dynamically adjusts weight distribution based on the characteristics of different datasets, enabling excellent performance across various scenarios.
IV Experiments
This section first introduces all 10 datasets used in the experiment, along with the quantitative and qualitative analyses of the experiment. Subsequently, we provide a detailed description of the parameter configurations used in the experiment. Finally, we outline the ablation study conducted for our proposed method.
IV-A Datasets
To validate the efficient segmentation of whole-body risk organs and lesions by MCFNet, we conduct experiments on ten datasets. The detailed descriptions of each dataset are as follows:
IV-A1 HaN (2015)
This is a public dataset for risk organs in the head and neck region, containing CT and MRI images of patients’ head and neck. These patients underwent CT scans and T1-weighted MR imaging to support image-guided radiotherapy planning. The segmentation labels provided by the dataset are stored separately; we preprocessed the segmentation labels of six regions, merging them into one segmentation label file and classifying them in the order of brainstem, mandible, left parotid gland, right parotid gland, left submandibular gland, and right submandibular gland. After preprocessing the dataset, we divide the image data of 17 tumor patients into 14 cases for the training set and 3 cases for the test set, with the training set containing 1924 slices.
IV-A2 SegTHOR (2019)
This dataset is a CT dataset specifically for the segmentation of thoracic risk organs (OARs) and was included in the ISBI 2019 challenge. In this dataset, the risk organs for the segmentation task include the heart, trachea, aorta, and esophagus, which differ in spatial and appearance characteristics. After preprocessing the dataset, we divide the image data of 40 cases into 32 cases for the training set and 8 cases for the test set, with the training set containing 6035 slices.
IV-A3 CHAOS (2019)
As a multi-modal abdominal medical image segmentation dataset, it provides paired multi-modal CT and MRI data along with corresponding annotation information. MRI imaging data is obtained from two different sequences (T1-DUAL and T2-SPIR). Since the CT images only provide liver annotations, and the MRI images provide annotations for four organs (liver, spleen, left kidney, and right kidney), we select MRI imaging of 20 patients with two modalities. After preprocessing the dataset, we divide the image data of 40 cases into 32 cases for the training set and 8 cases for the test set, with the training set containing 1026 slices.
IV-A4 LiTS (2017)
This dataset focuses on the segmentation of the liver and its tumors in CT images. The data were collected from seven different medical centers and were included in the ISBI 2017, MICCAI 2017, and MICCAI 2018 segmentation challenges. From the original dataset of 131 patients, we randomly select data from 40 patients. After preprocessing the images, we obtain 32 cases for the training set and 8 cases for the test set, with the training set containing 4599 slices.
IV-A5 KiTS (2019)
This dataset is a CT dataset for the segmentation of kidneys and their tumors. It was included in the MICCAI 2019 challenge, which aimed to accelerate the development of reliable semantic segmentation methods for kidneys and kidney tumors. All data were sourced from arterial phase abdominal CT scans of 300 kidney cancer patients who underwent partial or radical nephrectomy, with ground truth labels for kidneys and tumors provided. From the 210 publicly available cases, we randomly select data from 80 patients. After preprocessing the images, we obtain 64 cases for the training set and 16 cases for the test set, with the training set containing 4754 slices.
IV-A6 KiTS (2023)
This dataset focuses on the segmentation of kidneys, tumors, and cysts, and is an extension of the KiTS (2019) dataset. It includes additional segmentation labels for cysts and increases the total data volume from 210 cases to 489 cases. This dataset was included in the MICCAI 2023 challenge. From the 489 patients’ image data, we randomly select data from 80 patients for model validation. After preprocessing the data, we divide it into 64 cases for the training set and 16 cases for the test set, with the training set containing 4968 slices.
IV-A7 LUNG (2018)
This dataset is a subtask of the MSD Medical Image Segmentation Decathlon Challenge, aiming to segment lung tumors from CT images, highlighting the challenge of segmenting small targets within a large context. The official dataset includes thin-slice CT scans from 63 non-small cell lung cancer patients. After preprocessing the images, we divide the dataset into 50 cases for the training set and 13 cases for the test set, with the training set containing 1264 slices.
IV-A8 CPCGEA (2023)
This dataset includes MRI data of prostate cancer patients, comprising two modalities: DWI and T2WI sequences. The images suffer from poor clarity and additional noise due to environmental conditions or image sensor issues. Errors during image acquisition can also lead to artifacts such as streaks and shadows. The dataset consists of 172 cases with manually annotated prostate cancer regions. In our experiments, we randomly divide the data into 139 cases for the training set and 33 cases for the test set, with the training set containing 832 slices.
IV-A9 KNEE (2023)
This dataset consists of MRI images of patients’ leg bones, focusing on the segmentation of the femur. It includes MRI image data from 59 patients. After preprocessing the images, we obtain 47 cases for the training set and 12 cases for the test set, with the training set containing 8171 slices.
((a))
((b))
((c))
((d))
Figure 5: Training loss trends of the proposed method and seven comparative methods across four datasets.
IV-A10 AutoPET (2023)
AutoPET is a large-scale PET/CT dataset focusing on whole-body tumor segmentation, included in the MICCAI 2022 and MICCAI 2023 challenges. The official dataset comprises 1014 paired PET-CT images from 900 patients. Each data instance includes a three-dimensional whole-body FDG-PET image, a three-dimensional whole-body CT image, and a manually annotated three-dimensional tumor mask. Notably, all CT and PET images were acquired in the same scanning session using the same PET/CT scanner, ensuring high anatomical consistency. The dataset primarily covers patients with histologically confirmed malignant melanoma, lymphoma, and lung cancer, as well as negative control patients who underwent FDG-PET/CT scans at the University Hospital of Tübingen and LMU University Hospital in Munich, Germany. In our experiments, we select the PET images and annotations. We randomly divide the data from 80 patients into 64 cases for the training set and 16 cases for the test set, with the training set containing 2558 slices.
IV-B Evaluation Metrics
In all experiments, we comprehensively evaluate the performance of the segmentation models by applying appropriate metrics based on the dataset type.
For segmentation tasks on multi-class datasets, we use two key evaluation metrics: the Dice Similarity Coefficient (DSC) and the 95% Hausdorff Distance (HD95). DSC measures the overlap between the predicted segmentation and the ground truth, with higher values indicating better segmentation performance. HD95 assesses the boundary deviations in the segmentation results, where smaller values reflect higher accuracy in boundary prediction.
For segmentation tasks on single-class datasets, we use four evaluation metrics: DSC, HD95, Recall, and Precision. In addition to DSC and HD95, Recall measures the model’s ability to detect all instances of the target class, showing how well the model captures the entire target region. Precision measures the proportion of correctly identified target regions among all predicted positives, evaluating the model’s ability to avoid false positives. By combining these four metrics, we can more comprehensively and accurately assess the performance of single-class segmentation models.
Figure 6: Visual examples of segmentation on the HaN, SegTHOR, CHAOS, and LiTS datasets.
IV-C Implementation Detail
All experiments are conducted on a NVIDIA A6000 GPU. We standardize the resolution of all input images to $256\times 256$. The max training epoch is 300 with a batch size of 16 for all experiments. We use the Adam optimizer for model training with an initial learning rate of 0.001 and a cosine annealing learning rate decay strategy. The weight decay is set to 0.0001. The loss function used is a combination of Dice Loss and Cross-Entropy Loss. It can be defined as:
$\begin{split}L(Y,P)=&1-\sum_{i=1}^{I}(\lambda\frac{2*\sum_{n=1}^{N}{Y}_{n,i}\cdot{P}_{n,i}}{\sum_{n=1}^{N}{Y}_{n,i}^{2}+\sum_{n=1}^{N}{P}_{n,i}^{2}}\\ &+\sum_{n=1}^{N}{Y}_{n,i}log{P}_{n,i}),\end{split}$ (12)
$I$ is the number of classes, $N$is the total number of voxel. ${Y}_{n,i}$ and ${P}_{n,i}$ are the ground truth and output probability for the $i$-th class at voxel $n$, respectively.
IV-D Comparsion With Other Methods
We employ seven state-of-the-art methods to validate the performance of MCFNet: U-Net, TransUNet, Swin-Unet, UCTransNet, TransCASCADE, Cascaded MERIT, and FSCA-Net. We validate the segmentation performance of our proposed method against seven comparison models across all ten datasets. In Fig. 5, the training loss trends of all methods are illustrated across the four datasets. Fig. 5 shows visual examples of segmentation on all datasets and we highlight incorrectly segmented areas with red boxes. The specific qualitative and quantitative analysis results are as follows:
IV-D1 Experiments on HaN
Our proposed method and seven comparison models segment six regions in this dataset. As shown in the Table I, the Cascaded MERIT model achieves excellent segmentation performance, with an average DSC of 74.71% and an average HD95 of 2.532mm. Our designed MCFNet improves the average DSC by 2.81% and reduces the average HD95 by 0.079mm compared to Cascaded MERIT. The experimental results in Table I show that our model significantly improves the segmentation of the left submandibular gland.
As shown in the first row of Fig. 6, we observe that our proposed method does not exhibit under-segmentation or over-segmentation compared to the comparison models. For example, UNet incorrectly identifies the mandibular region, while Swin-Unet fails to fully recognize the brainstem region, resulting in under-segmentation. Additionally, UCTransNet fails to accurately distinguish between the left and right parotid glands.
TABLE I: COMPARISONS WITH STATE-OF-THE-ART MODELS ON THE HAN DATASET.
[…]
UNet 72.98 3.260 77.44 89.45 73.68 76.09 59.08 62.17
TransUNet 71.48 5.412 76.12 90.81 71.97 71.16 56.96 61.82
Swin-Unet 68.75 7.541 73.51 91.67 70.53 73.27 50.63 52.85
UCTransNet 70.36 4.328 77.26 88.85 70.38 71.03 61.23 53.43
TransCASCADE 71.56 3.184 79.39 87.67 70.75 75.20 51.44 64.89
Cascaded MERIT 74.71 2.532 79.17 92.14 75.70 78.57 54.78 67.89
FSCA-Net 73.31 4.822 75.52 90.19 73.62 72.36 64.06 64.11
Ours 77.52 2.453 79.52 93.47 77.59 78.61 68.02 67.92
TABLE II: COMPARISONS WITH STATE-OF-THE-ART MODELS ON THE SEGTHOR DATASET.
[…]
UNet 75.30 6.190 56.06 84.27 82.25 78.62
TransUNet 80.79 4.542 64.24 88.31 84.70 85.91
Swin-Unet 71.32 7.208 47.12 85.60 77.43 75.15
UCTransNet 80.86 5.115 64.36 87.91 85.32 85.85
TransCASCADE 77.18 4.608 58.54 89.26 80.37 80.55
Cascaded MERIT 78.92 4.272 59.16 89.30 83.65 83.58
FSCA-Net 80.75 4.947 65.56 85.84 84.48 87.12
Ours 83.27 4.187 69.37 89.55 87.75 87.40
TABLE III: COMPARISONS WITH STATE-OF-THE-ART MODELS ON THE CHAOS DATASET.
[…]
UNet 90.82 8.990 94.02 90.42 90.43 88.43
TransUNet 91.19 5.902 94.40 91.92 91.06 87.40
Swin-Unet 90.17 4.366 93.61 90.23 89.65 87.20
UCTransNet 88.94 6.522 91.26 90.05 86.96 87.49
TransCASCADE 90.26 3.695 93.66 91.61 89.79 85.99
Cascaded MERIT 90.69 3.914 93.57 91.25 91.74 86.18
FSCA-Net 89.83 9.277 93.55 91.12 88.25 86.41
Ours 92.80 3.676 94.87 93.45 92.05 90.83
IV-D2 Experiments on SegTHOR
Our proposed method and seven comparison models segment the trachea, heart, aorta, and esophagus in this dataset. According to the experimental results in Table II, current state-of-the-art methods such as TransUNet, UCTransNet, and FSCA-Net achieve over 80% in average DSC and reduce the average HD95 to below 5mm. Our proposed method achieves an average DSC of 83.27% and an average HD95 of 4.187mm, with MCFNet obtaining the best performance in segmenting the trachea, heart, aorta, and esophagus.
As shown in the second row of Fig. 6, our designed model accurately identifies and segments the risk organs. However, UNet, TransUNet, UCTransNet, Cascaded MERIT, and FSCA-Net all mistakenly identify the background as the aorta. Additionally, Swin-Unet incorrectly identifies the esophagus region as the heart.
IV-D3 Experiments on CHAOS
The comparison results are shown in Table III. We can see that the cascade networks TransCASCADE [33] and Cascaded MERIT achieve relatively good segmentation results for the four abdominal organs, with average DSCs of 90.26% and 90.69%, and average HD95s of 3.695mm and 3.914mm, respectively. The traditional TransUNet achieves an average DSC of 91.19%, making it the best-performing model among the comparisons. Our designed model achieves an average DSC of 92.80% and reduces the average HD95 to 3.676mm, significantly outperforming all comparison models.
As shown in the third row of Fig. 6, most comparison models exhibit under-segmentation for the liver region, and Cascaded MERIT also shows ambiguous identification for the spleen and right kidney. Our designed model generally achieves accurate segmentation, but there is still room for improvement in the segmentation of the spleen’s edges.
TABLE IV: COMPARISONS WITH STATE-OF-THE-ART MODELS ON THE LITS DATASET.
[…]
UNet 64.10 31.754 92.11 36.11
TransUNet 71.46 25.951 90.97 51.93
Swin-Unet 66.09 30.759 91.83 40.35
UCTransNet 66.77 29.473 92.66 40.87
TransCASCADE 72.56 25.351 94.16 50.96
Cascaded MERIT 72.36 24.521 93.25 51.47
FSCA-Net 68.54 23.897 93.04 44.04
Ours 73.79 23.326 94.22 53.35
TABLE V: COMPARISONS WITH STATE-OF-THE-ART MODELS ON THE KITS (2019) DATASET.
[…]
UNet 65.20 23.039 87.99 42.41
TransUNet 66.41 34.096 89.52 43.31
Swin-Unet 65.21 33.455 85.09 45.33
UCTransNet 64.31 43.377 87.25 41.37
TransCASCADE 68.14 25.023 89.78 46.43
Cascaded MERIT 68.50 22.970 90.29 46.38
FSCA-Net 60.59 41.617 85.97 35.22
Ours 70.71 20.522 90.50 50.92
IV-D4 Experiments on LiTS
All results for this dataset are shown in Table IV. We can see that TransCASCADE is the best-performing advanced segmentation method, achieving an average DSC of 72.56%. For liver tumor segmentation, TransUNet achieves a DSC of 51.93%, which is currently the best. Our designed model achieves better segmentation results for tumors, with an average DSC of 73.79% and a DSC of 53.35% for tumor regions.
As shown in the fourth row of Fig. 6, our designed model, like many comparison models, achieves relatively accurate segmentation for the liver and liver tumors. However, there is a gap in the detailed segmentation of the tumor region edges. Additionally, models like UNet, due to the small size of the segmentation target, may fail to recognize the tumor region.
Figure 7: Visual examples of segmentation on the KiTS (2019), KiTS (2023), LUNG, CPCGEA and KNEE datasets.
TABLE VI: COMPARISONS WITH STATE-OF-THE-ART MODELS ON THE KITS (2023) DATASET.
[…]
UNet 51.79 30.987 90.14 33.02 32.20
TransUNet 58.31 29.390 90.94 42.31 41.68
Swin-Unet 49.41 31.283 87.02 35.52 25.69
UCTransNet 47.19 44.911 87.92 18.86 34.79
TransCASCADE 59.38 26.173 91.21 44.76 42.17
Cascaded MERIT 61.01 23.033 91.29 45.05 46.69
FSCA-Net 46.55 28.441 90.67 21.83 27.16
Ours 64.07 22.770 91.74 45.18 55.31
IV-D5 Experiments on KiTS (2019)
This dataset focuses on the segmentation of kidneys and kidney tumors. As shown in Table V, our designed model achieves an average DSC that is 2.21% higher than the state-of-the-art model Cascaded MERIT and an average HD95 that is 2.448mm lower. Additionally, MCFNet improves the tumor segmentation DSC by 4.49% compared to TransCASCADE.
As shown in the fifth row of Fig. 7, we demonstrate the segmentation effects of the models on an image containing both the kidney and tumor. Swin-Unet exhibits scattered segmentation in the tumor region, with dispersed pixels. UCTransNet mistakenly identifies some background areas as tumors. Even more problematic, FSCA-Net fails to recognize the tumor region altogether.
TABLE VII: COMPARISONS WITH STATE-OF-THE-ART MODELS ON THE LUNG DATASET.
[…]
(%, mean) (mm, mean) (%, mean) (%, mean)
U-Net 70.50 37.913 66.07 80.04
TransUNet 70.72 25.218 60.56 82.06
SwinUnet 63.41 25.131 57.47 77.29
UCTransNet 70.75 32.384 65.90 80.38
TransCASCADE 68.31 28.551 65.01 76.51
Cascaded MERIT 70.43 28.554 67.33 83.34
FSCA-Net 72.10 25.545 68.07 82.23
Ours 74.49 23.505 71.06 83.52
Figure 8: Visual examples of segmentation on the AutoPET dataset.
IV-D6 Experiments on KiTS (2023)
This dataset builds upon the segmentation labels of kidneys and kidney tumors from KiTS (2019) by adding segmentation for cysts. As shown in Table VI, traditional networks do not perform well in the segmentation of tumors and cysts. For example, the best-performing TransUNet achieves an average DSC of only 42.31% for tumors and 42.68% for cysts. In contrast, our designed model improves the DSC for cyst segmentation to 55.31%, and MCFNet achieves the best performance in segmentation for kidneys, kidney tumors, and cysts.
TABLE VIII: COMPARISONS WITH STATE-OF-THE-ART MODELS ON THE CPCGEA DATASET.
[…]
(%, mean) (mm, mean) (%, mean) (%, mean)
U-Net 64.05 9.312 70.34 65.99
TransUNet 63.18 8.517 70.02 66.14
SwinUnet 60.94 7.185 66.20 65.24
UCTransNet 60.25 11.374 67.59 64.78
TransCASCADE 65.44 6.646 68.85 69.14
Cascaded MERIT 65.50 7.738 70.47 68.57
FSCA-Net 65.24 6.792 68.88 68.37
Ours 67.28 6.064 71.48 69.43
TABLE IX: COMPARISONS WITH STATE-OF-THE-ART MODELS ON THE KNEE DATASET.
[…]
(%, mean) (mm, mean) (%, mean) (%, mean)
U-Net 98.99 1.089 98.65 99.23
TransUNet 98.92 1.094 98.74 99.12
SwinUnet 98.25 1.228 97.79 98.72
UCTransNet 99.08 1.055 98.82 99.01
TransCASCADE 99.02 1.000 98.89 99.16
Cascaded MERIT 99.09 1.000 98.95 99.24
FSCA-Net 97.59 5.362 97.00 98.43
Ours 99.14 1.000 98.99 99.29
TABLE X: COMPARISONS WITH STATE-OF-THE-ART MODELS ON THE AUTOPET DATASET.
[…]
(%, mean) (mm, mean) (%, mean) (%, mean)
U-Net 58.49 30.832 59.82 68.62
TransUNet 61.95 22.329 60.22 72.98
SwinUnet 50.55 31.925 53.67 57.58
UCTransNet 58.05 28.279 59.77 65.51
TransCASCADE 55.18 24.160 49.92 73.31
Cascaded MERIT 58.32 23.407 57.10 69.09
FSCA-Net 61.53 26.707 62.07 70.76
Ours 63.87 22.251 65.35 73.91
As shown in the sixth row of Fig. 7, the addition of cyst segmentation labels increases the difficulty of segmenting small target areas. Models like UNet, TransUNet, and Swin-Unet demonstrate shortcomings in segmenting cysts effectively. Additionally, while UCTransNet accurately identifies cyst regions, it struggles to maintain accurate recognition of tumor regions simultaneously.
IV-D7 Experiments on LUNG
In this dataset, we select four evaluation metrics to validate the model’s segmentation performance, as shown in Table VII. It is observed that the advanced model FSCA-Net achieves the best results in terms of DSC and Recall, while Cascaded MERIT achieves the best Precision among existing models, reaching 82.23%. Our designed model achieves the best performance across DSC, HD95, Recall, and Precision metrics. Specifically, MCFNet improves DSC by a significant 2.39% and Recall by 2.99% compared to FSCA-Net.
As shown in the seventh row of Fig. 7, for single-label segmentation targets, most models correctly identify tumor regions except for Swin-Unet. UNet identifies tumor regions but also mistakenly includes background areas as tumors. Swin-Unet not only fails to accurately identify tumor regions but also misidentifies vascular regions in the background as tumors.
IV-D8 Experiments on CPCGEA
As shown in Table VIII, we compare seven comparison models for the segmentation performance of prostate cancer. Since we are dealing with single-label image segmentation, we use four evaluation metrics to comprehensively assess model performance. Our designed model improves upon the current best model by 1.78% in DSC, reduces HD95 by 0.582mm, increases Recall by 1.01%, and improves Precision by 0.29%.
As shown in the eighth row of Fig. 7, we continue the work initiated by our previously proposed model, FSCA-Net, and conducted further exploration on CPCGEA. We add experiments with TransCASCADE and Cascaded MERIT, revealing that they exhibit unclear boundary segmentation for prostate cancer regions and struggle to delineate the edges of tumor areas effectively.
TABLE XI: ABLATION EXPERIMENTS ON FIVE DATASETS.
Method SegTHOR CHAOS LiTS KiTS19 CPCGEA
[…]
(%, mean) (mm, mean) (%, mean) (mm, mean) (%, mean) (mm, mean) (%, mean) (mm, mean) (%, mean) (mm, mean) (%, mean) (%, mean)
SEB 78.74 5.063 90.17 6.051 66.47 26.199 67.32 33.096 63.18 8.520 70.02 66.14
FCB 80.75 4.947 89.83 9.277 68.54 23.897 60.59 41.617 65.24 6.792 68.88 68.37
SEB+Adaptive-MFA 79.72 5.220 90.44 4.366 67.75 35.554 67.84 30.069 63.75 7.746 75.76 59.78
SCB+Adaptive-MFA 81.43 5.040 90.01 8.253 70.56 25.951 64.31 43.377 65.44 6.646 68.85 69.14
SEB+FCB 82.87 5.306 91.91 6.001 73.25 26.803 68.32 23.997 65.50 7.738 70.47 68.57
SEB+FCB+Adaptive-MFA 83.27 4.187 92.80 3.676 73.79 23.326 70.71 20.522 67.28 6.064 71.48 69.43
Figure 9: The experiments shown in the figure are conducted on the CPCGEA dataset. (a) Visualization of model performance and parameter size. (b) Visualization of model performance and FLOPs.
IV-D9 Experiments on KNEE
As shown in Table IX, all models achieve an average DSC for femur segmentation exceeding 97.5%, with our designed model achieving the highest at 99.14%. Additionally, HD95, Recall, and Precision also achieve optimal performance, although the improvement in segmentation performance compared to advanced models is relatively limited.
As depicted in the ninth row of Fig. 7, we can see that all models accurately segment the femur region. However, Swin-Unet and UCTransNet exhibit minor segmentation errors in very small areas, although the overall differences are minimal.
IV-D10 Experiments on AutoPET
For whole-body tumor segmentation, our designed model and seven comparison models are evaluated using four metrics, as shown in the Table X. Among the current advanced methods, TransUNet achieves the best DSC of 61.95%, the best HD95 of 22.329mm, and the best Precision of 72.98%. Additionally, FSCA-Net achieves a Recall of 62.07%. However, our designed model MCFNet improves by 1.92% in DSC, reduces HD95 by 0.078mm, increases Recall by 3.28%, and improves Precision by 0.93%.
As shown in Fig. 8, four segmentation cases from the AutoPET dataset are presented. The erroneous segmentation areas are all marked with red boxes. In the whole-body tumor segmentation images, there are many background areas that appear white and have smooth edges but are not tumors, leading the segmentation models to mistakenly identify tumor regions. From a comprehensive observation of the four cases, it can be seen that most models exhibit confusion between the background and tumor regions. Specifically, UNet and TransUNet have issues with over-segmentation and under-segmentation. Swin-Unet, UCTransNet, and FSCA-Net all display errors in segmentation, while TransCASCADE and Cascaded MERIT show cases of under-segmentation. Although the MCFNet model has performed relatively well in tumor segmentation, there is still room for improvement in segmenting the detailed boundaries of tumor regions.
IV-E Ablation Study
To validate the effectiveness of our proposed module, we conduct experiments on five datasets and use multiple evaluation metrics to comprehensively assess the module’s effectiveness. As shown in Table XI, after applying the adaptive loss aggregation strategy to SEB and FCB, the segmentation performance significantly improves across all datasets. Additionally, compared to using SEB and FCB individually, cascading SEB and FCB enhances the model’s segmentation performance, and further applying the adaptive loss aggregation strategy on top of this results in the best segmentation performance.
IV-F Discussion
MCFNet integrates low parameter count, low computational cost, and high segmentation performance. Our designed model effectively balances performance with parameter and computational cost, demonstrating significant advantages over existing state-of-the-art methods. To illustrate this, we conduct experiments on the CPCGEA dataset and compare our model’s results with several benchmark models. These experimental results are visualized in figures, clearly showing the performance of the MCFNet model. As shown in Fig. 9, our model not only exhibits excellent segmentation performance but also achieves or surpasses the performance of existing methods while maintaining a smaller parameter size and lower computational cost. This indicates that our model offers higher efficiency and better scalability in practical applications. The optimization of performance, computational cost, and parameter size makes our model suitable for resource-constrained environments and lays a solid foundation for its future promotion and application in various scenarios.
V Conclusion
We design a novel multi-scale, multi-resolution cascade fusion network called MCFNet. By capturing rich image features from multi-scale and multi-resolution images, our designed model is better suited to various application scenarios. We propose an adaptive loss aggregation strategy called Adaptive-MFA to optimize model training and performance. We conduct experiments on multiple datasets covering whole-body risk organs and tumors to fully validate the robustness and generalization capability of the model. Experimental results show that our MCFNet exhibits excellent segmentation performance and good generalization ability across multiple datasets.
In practical clinical applications, the efficient segmentation capability of MCFNet can significantly enhance the accuracy and efficiency of medical image analysis, reduce the workload of doctors, and improve the precision of diagnosis and treatment. By accurately segmenting risk organs and tumor regions, MCFNet can assist doctors in better formulating treatment plans, thereby improving patient outcomes. Additionally, the good generalization ability of MCFNet means it can adapt to image data generated by different hospitals and equipment, making it a promising tool for widespread use in clinical practice. This can drive advancements and progress in the field of medical image analysis.

### v2

@@
-Multi-scale Cascaded Large-Model for Whole-body ROI Segmentation
+Multi-scale Cascaded Foundation Model for Whole-body Organs-at-risk Segmentation
 Abstract
-Organs-at-risk segmentation is critical for ensuring the safety and precision of radiotherapy and surgical procedures.
-However, existing methods for organs-at-risk image segmentation often suffer from uncertainties and biases in target selection, as well as insufficient model validation experiments, limiting their generality and reliability in practical applications.
-To address these issues, we propose an innovative cascaded network architecture called the Multi-scale Cascaded Fusing Network (MCFNet), which effectively captures complex multi-scale and multi-resolution features.
-MCFNet includes a Sharp Extraction Backbone and a Flexible Connection Backbone, which respectively enhance feature extraction in the downsampling and skip-connection stages.
-This design not only improves segmentation accuracy but also ensures computational efficiency, enabling precise detail capture even in low-resolution images.
-We conduct experiments using the A6000 GPU on diverse datasets from 671 patients, including 36,131 image-mask pairs across 10 different datasets.
-MCFNet demonstrates strong robustness, performing consistently well across 10 datasets.
-Additionally, MCFNet exhibits excellent generalizability, maintaining high accuracy in different clinical scenarios.
-We also introduce an adaptive loss aggregation strategy to further optimize the model training process, improving both segmentation accuracy and efficiency.
-Through extensive validation, MCFNet demonstrates superior performance compared to existing methods, providing more reliable image-guided support.
-Our solution aims to significantly improve the precision and safety of radiotherapy and surgical procedures, advancing personalized treatment.
+Accurate segmentation of organs-at-risk (OARs) is vital for safe and precise radiotherapy and surgery.
+Most existing studies segment only a limited set of organs or regions, lacking a systematic treatment of OARs segmentation.
+We present a Multi-scale Cascaded Fusion Network (MCFNet) that aggregates features across multiple scales and resolutions.
+MCFNet consists of a Sharp Extraction Backbone for the downsampling path and a Flexible Connection Backbone for skip-connection fusion, strengthening representation learning in both stages.
+This design improves boundary localization and preserves fine structures while maintaining computational efficiency, enabling reliable performance even on low-resolution inputs.
+Experiments on an NVIDIA A6000 GPU using 36,131 image–mask pairs from 671 patients across 10 datasets show consistent robustness and strong cross-dataset generalization.
+An adaptive loss-aggregation strategy further stabilizes optimization and yields additional gains in accuracy and training efficiency.
+Through extensive validation, MCFNet outperforms existing methods, excelling in organ segmentation and providing reliable image-guided support for computer-aided diagnosis.
+Our solution aims to improve the precision and safety of radiotherapy and surgery while supporting personalized treatment, advancing modern medical technology.
 The code has been made available on GitHub: https://github.com/Henry991115/MCFNet.
-Index Terms:
-Whole-body ROI segmentation, multi-scale cascaded Large-Model, linear attention transformer, adaptive loss aggregation strategy.
 I Introduction
-Medical image segmentation aims to quickly and accurately separate different tissues, organs, or lesion areas, facilitating clinical applications such as disease analysis, lesion detection, and treatment planning [1, 2].
-In the field of medical image analysis, accurately identifying and segmenting malignant tumors is crucial for early cancer diagnosis, treatment planning, and efficacy evaluation.
-However, malignant tumors are often complex in shape, vary in size, have blurred boundaries, and are closely connected to surrounding tissues, posing significant challenges for accurate tumor identification and segmentation [3, 4, 5].
-Accurately identifying and locating organs-at-risk (OAR) can help doctors and medical teams plan and execute treatment plans.
-In radiation therapy, precise identification of OAR can assist doctors in minimizing the impact of treatment on surrounding healthy tissues, thereby reducing adverse effects caused by the treatment [6, 7].
-Tumor and OAR segmentation are fundamental to treatment planning in radiation therapy.
-Tumor segmentation helps doctors determine the target area for accurately aiming the radiation dose.
-Meanwhile, OAR segmentation allows doctors to identify and protect surrounding healthy tissues, avoiding unnecessary damage during treatment [8, 9].
-With the emergence of deep learning models, traditional methods constrained by human, financial, and time costs are gradually being abandoned [10, 11, 12, 13].
-Deep learning-based models can automatically learn features from images and efficiently extract features of complex structures and textures, enabling the identification of ROI contours and facilitating subsequent medical image analysis [14, 15].
-Most current efficient segmentation models require a substantial amount of high-performance GPU computing resources for training and have a strong dependency on large-scale annotated data.
-However, obtaining high-quality annotated data requires a significant amount of time and human resources [16, 17].
-Most current deep learning models for segmentation are based on UNet [18] and Transformer [19] architectures.
-UNet is renowned for its efficient and simple structure, which can effectively capture local features in images, while Transformers are known for their outstanding ability to capture global information [20, 21, 22].
-The combination of UNet’s local feature extraction capabilities with the Transformer’s global information capture ability allows segmentation models to consider both local details and overall structure in images, thereby improving segmentation accuracy and robustness.
-This combination provides a more comprehensive and efficient solution for medical image analysis, offering strong support for advancements in the medical imaging field.
-However, the segmentation performance of most current methods is limited by the application scenario, and the robustness and generalization abilities of the models are usually weak.
-High-performance models like TransUNet [23], Swin-Unet [24], and MISSFormer [25] have demonstrated their capabilities primarily in segmenting eight organs in the heart and abdominal regions.
-UCTransNet [26] has expanded its tasks to include the segmentation of glands and cell nuclei.
-Nevertheless, in our actual experiments, we found that the segmentation performance of these methods varies significantly when applied to datasets with different image characteristics.
-Existing works primarily focus on segmenting certain organs-at-risk and lesions, but they cannot segment multiple regions of the entire body, highlighting the limited application scenarios and the relatively poor robustness and generalization capabilities of these models.
-To address this, we propose conducting experiments on datasets that cover multiple organs-at-risk and tumors across the entire body.
-To enhance the model’s generalization ability for different image characteristics, we intend to capture image features from multiple angles using multi-scale and multi-resolution images, enabling the model to better adapt to current application scenarios.
-Finally, we design the FCB (Flexible Connection Backbone), which is targeted for multi-region segmentation, as one of the strong backbones.
-Additionally, we design a SEB (Sharp Extraction Backbone) to assist the FCB in better feature extraction.
-Cascading the two backbones can ensure high segmentation performance while designing the model to enhance its robustness and generalization ability.
-Our contributions are summarized as follows:
+Medical image segmentation aims to accurately separate different tissues, organs, or lesion areas, supporting clinical tasks like disease analysis, lesion detection, and treatment planning [1, 2].
+Accurate segmentation of malignant tumors is crucial for early cancer diagnosis, treatment planning, and efficacy evaluation.
+However, tumors often have complex shapes, blurred boundaries, and are closely connected to surrounding tissues, posing challenges for segmentation [3, 4, 5, 6].
+Identifying and locating OARs is vital in radiation therapy, helping minimize damage to healthy tissues and improving treatment outcomes [7, 8, 9].
+Tumor and OAR segmentation are essential for precise treatment planning, enabling targeted radiation while protecting healthy tissues [10, 11].
+Figure 1: MCFNet is trained on ten diverse datasets to perform whole-body organs-at-risk (OARs) segmentation, covering the head & neck, thorax, abdomen, prostate, and femur regions.
+With the rise of deep learning, traditional methods constrained by human and financial costs are being replaced [12, 13, 14, 15].
+Deep learning models can automatically extract complex features from images, enabling efficient OARs contour identification and further medical analysis [16, 17, 18].
+However, most current segmentation models require substantial GPU resources for training and large-scale annotated datasets, which are time-consuming and resource-intensive to create [19, 20].
+Most current deep learning segmentation models are based on UNet [21] and Transformer [22] architectures.
+UNet is known for its simple structure that effectively captures local features, while Transformers excel at capturing global information [23, 24, 25].
+Combining UNet’s local feature extraction with the Transformer’s global information capture improves segmentation accuracy and robustness by considering both local details and the overall structure in images.
+This combination enhances medical image analysis, supporting progress in the field [26, 27, 28].
+However, the segmentation performance of existing methods is often limited by the application scenario, with weak robustness and generalization abilities.
+High-performance models such as TransUNet [29], Swin-Unet [30], and MISSFormer [31] perform well in segmenting organs in the heart and abdominal regions, while UCTransNet [32] can segment glands and cell nuclei.
+Nevertheless, in our actual experiments, we find that the segmentation performance of these methods varies significantly when applied to datasets with different image characteristics [33].
+In this paper, we propose MCFNet, a multi-scale cascaded fusion framework for robust segmentation of OARs and lesions under diverse imaging protocols and anatomical regions.
+MCFNet synchronizes multi-scale and multi-resolution inputs through two complementary branches: the Flexible Connection Backbone (FCB) for coarse global semantics and the Sharp Extraction Backbone (SEB) for fine boundary details.
+A Linear Attention Transformer is embedded in FCB to model long-range dependencies efficiently, while SEB focuses on high-frequency structures.
+An adaptive multi-scale loss aggregation strategy further enhances accuracy and training stability.
+As shown in Fig.
+1, we evaluate MCFNet on ten datasets spanning head-neck, thorax, abdomen, prostate, and femur regions.
+Extensive experiments and ablation studies demonstrate its strong robustness, generalization, and parameter efficiency.
+The main contributions of our work are as follows:
   • $\bullet$
-    Novel Multi-scale Cascaded Large-Model: We propose a novel Multi-scale Cascaded U-shaped Network for whole-body ROI segmentation.
-This network cascades two backbone networks, namely our designed FCB (Flexible Connection Backbone) and SEB (Sharp Extraction Backbone).
-The two backbones are used to handle images of different scales and resolutions, with feature fusion occurring in the four skip connections and the down-sampling bottom layers of the network.
+    We propose MCFNet, a novel multi-scale cascaded network that integrates the SEB and FCB to encode and fuse features across scales and resolutions.
+Feature fusion is performed at four skip connections and the bottleneck layers, enabling both boundary fidelity and semantic consistency.
+Comprehensive evaluations on 10 heterogeneous datasets show that the proposed framework is robust, computationally efficient, and generalizes well.
   • $\bullet$
-    Innovative Adaptive Loss Aggregation Strategy: We introduce a novel adaptive loss aggregation strategy named Adaptive-MFA (Adaptive Multi-scale Feature-Mixing Loss Aggregation).
-By using this strategy, we can create new prediction maps by combining available prediction maps, thereby optimizing the model training as much as possible.
+    We develop a Linear Attention Transformer Block (LAT) within the FCB to enhance feature extraction along the channel and spatial axes.
+LAT reduces quadratic complexity to linear, thereby accelerating convergence and improving scalability.
+This design enables the FCB to capture global dependencies with fewer parameters, achieving a favorable accuracy–efficiency trade-off.
   • $\bullet$
-    Extensive Validation for Robustness and Generalization: To thoroughly validate the model’s robustness and generalization capabilities, we selected ten datasets targeting organs-at-risk and tumors throughout the body for experiments.
-The segmented regions include the brain, thoracic cavity, abdomen, and lower limbs, with two of these datasets provided by our collaborating hospital.
+    We propose an Adaptive Multi-scale Feature-Mixing Loss Aggregation (Adaptive-MFA) strategy, which dynamically generates new prediction maps by combining multi-scale outputs.
+This adaptive aggregation stabilizes training and further improves segmentation performance.
+The structure of this study is as follows: Section II discusses related work.
+The methodology of our proposed model is detailed in Section III.
+The experiment configuration is introduced in Section IV.
+Section IV describes the results and discussion, and the effectiveness of the proposed method is validated.
+Finally, Section V presents the summarization of our research conclusions.
+Figure 2: Discussion on Model Complexity on the CPCGEA Dataset. (a) Visualization of model performance and parameter size. (b) Visualization of model performance and FLOPs.
 II Related Work
-In this section, we introduce commonly used architectures for medical image segmentation as well as cascade network architectures.
+This section introduces common architectures for medical image segmentation and cascade networks.
 II-A Medical Image Segmentation
-Medical image segmentation refers to accurately separating different structures (such as organs, lesion regions, etc.) in medical images, which is a crucial task in medical image analysis and computer-aided diagnosis.
-To address this task, research on U-shaped networks based on encoder-decoder structures has seen explosive growth.
-UNet innovatively introduced the encoder-decoder architecture, enabling the aggregation of features from different stages through skip connections.
-Subsequently, UNet++ [28] and UNet 3+ [29] improved the skip connection mechanism, further enhancing the effectiveness of feature aggregation.
-In recent years, the application of Transformers in medical image segmentation has attracted widespread attention.
-Originally used in the field of natural language processing, Transformers have shown great potential in computer vision tasks due to their powerful modeling capabilities and flexibility.
-Compared to traditional convolutional neural networks (CNNs), Transformer architectures primarily rely on self-attention mechanisms, allowing them to capture long-range dependencies in images globally.
-ViT [30] is one of the first models to apply Transformers to image classification tasks, processing images by splitting them into fixed-size patches and treating these patches as sequences inputted to the Transformer for processing.
-The same concept has been introduced into medical image segmentation, such as the TransUNet model, which significantly improves segmentation performance by combining the local feature extraction capability of UNet with the global contextual modeling capability of Transformers.
-Swin Transformer is a hierarchical Transformer architecture that further enhances computational efficiency and feature extraction capability through a sliding window mechanism.
-Swin-Unet applies this architecture to medical image segmentation tasks and achieves excellent results.
-Recently, Rahman et al.
-proposed the Cascaded MERIT [31] model, which effectively improves the accuracy and stability of medical image segmentation through techniques such as multiscale hierarchical feature extraction and cascaded attention decoding.
+Medical image segmentation, which involves accurately separating different structures like organs and lesions, is essential for medical image analysis and diagnosis.
+The introduction of encoder-decoder-based U-shaped networks, such as UNet, has significantly advanced this task, with subsequent improvements like UNet++ [34] and UNet 3+ [35] enhancing skip connection mechanisms for better feature aggregation.
+In recent years, Transformers have gained attention in medical image segmentation.
+Although initially designed for NLP, Transformers excel in computer vision tasks due to their powerful modeling and flexibility [36].
+Unlike CNNs, Transformer models employ self-attention to model long-range dependencies on a global scale.
+ViT [37] introduce Transformers to image classification, using fixed-size patches to segment images, treating them as sequences for processing.
+This idea has been adapted to medical image segmentation, exemplified by the TransUNet model, which combines UNet’s local feature extraction with the global context modeling of Transformers.
+The Swin Transformer, a hierarchical model with sliding windows, further boosts feature extraction and efficiency.
+The Cascaded MERIT model by Rahman et al.
+[38] has been shown to improve segmentation accuracy and stability using multiscale features and cascaded attention decoding.
+Figure 3: Illustration of the overall MCFNet architecture.
+The network is built upon two complementary backbones: the Sharp Extraction Backbone for fine details and the Flexible Connection Backbone for global semantics, which are integrated via a Cascaded Skip-connection Module and an Aggregation Module.
 II-B Cascaded Networks for Medical Image Segmentation
-In recent years, cascade networks have gained considerable attention in medical image segmentation.
-These networks concatenate multiple modules, each handling specific subtasks, gradually improving overall segmentation performance.
-Cascade networks can integrate feature information at different scales and decompose tasks into multiple stages for optimization, with each stage utilizing feature representations from the preceding stage [31].
-Information propagation is achieved through mechanisms such as skip connections and attention mechanisms.
-In experimental validation, cascade networks have demonstrated effectiveness and superiority in medical image segmentation tasks, offering significant potential as diagnostic and therapeutic tools in clinical medicine.
-Recently, the CASCADE model introduced a novel attention-based decoder that learns local (contextual) relationships between pixels, addressing the shortcomings of Transformer models in learning local relationships.
-G-CASCADE [32] refines multi-stage feature maps generated by a hierarchical Transformer encoder through graph convolutional blocks, contributing to improved accuracy and robustness of segmentation results through progressive refinement.
-The Cascaded MERIT [31] model improves generalization capability by introducing a multiscale self-attention mechanism.
-Traditional single-scale self-attention mechanisms may not effectively capture multiscale information in images, whereas Cascaded MERIT computes self-attention at multiple scales to comprehensively understand image features, thereby enhancing model performance.
-MCANet proposes a multi-scale cross-axis attention mechanism that efficiently captures information from different sizes and shapes of lesion regions or organs in medical images.
-By computing dual cross attentions between two parallel axial attentions, better global information capture is achieved, facilitating accurate segmentation of target areas in medical images.
+In recent years, cascade networks have become prominent in medical image segmentation.
+These networks concatenate multiple modules to handle specific subtasks, improving segmentation performance gradually.
+By integrating feature information at different scales and decomposing tasks into multiple stages, cascade networks optimize segmentation through mechanisms like skip connections and attention mechanisms [39].
+Experimental results show that cascade networks are effective and often superior for medical image segmentation, with consistent gains in Dice and HD95 across diverse benchmarks (e.g., cardiac structures, and tumor/OARs delineation)[40].
+Beyond accuracy, cascaded designs improve robustness to domain shifts and small, low-contrast targets, translating into more reliable auto-contouring and reduced manual correction time in clinical workflows—thereby offering tangible benefits for diagnostics, radiotherapy planning, and longitudinal treatment assessment[41, 42].
+The CASCADE model introduces an attention-based decoder that captures local pixel-wise dependencies, addressing a known limitation of Transformer architectures.
+G-CASCADE [43] employs graph-convolution blocks to refine hierarchical Transformer encoder features, improving segmentation accuracy and robustness.
+The Cascaded MERIT [38] model improves generalization by employing a multiscale self-attention mechanism, which better captures multiscale information than traditional single-scale attention mechanisms.
+MCANet [44] employs multi-scale cross-axis attention to model lesions and organs with heterogeneous sizes and shapes, strengthening global context and enabling precise target segmentation.
+Distinguished from existing studies, our method offers enhanced multi-scale modeling, demonstrating leading performance and generalization in robust multi-dataset evaluations.
+Figure 4: (a) Illustration of the Cascaded Skip-connection Module (CSM). (b) Illustration of the Linear Attention Transformer Block (LAT) in the CSM.
 III Proposed Method
-In this section, we introduce the multi-scale cascade network model MCFNet designed for whole-body ROI segmentation.
+In this section, we introduce the multi-scale cascade network model MCFNet designed for whole-body OARs segmentation.
 First, we provide an overall introduction to the model, then we describe each component of the model in detail, and finally, we explain the adaptive loss aggregation strategy we designed.
-Figure 1: Illustration of the overall architecture of MCFNet.
-III-A MCFNet for Whole-body ROI Segmentation
+III-A MCFNet for Whole-body OARs Segmentation
 Fig.
@@
 MCFNet consists of three parts: Cascaded Backbones, Decoders with cascaded skip connections, and Aggregation.
-The Cascaded Backbones module is composed of SEB (Acute Extraction Backbone) and FCB (Elegant Connection Backbone), enabling multi-scale and multi-resolution input for the model.
-The Decoders with cascaded skip connections module mainly achieves feature aggregation at skip connections and bridge layers in the cascade network, enhancing feature representation by fully utilizing feature information from different levels and different sub-networks.
+The Cascaded Backbones module is composed of Sharp Extraction Backbone (SEB) and Flexible Connection Backbone (FCB), enabling multi-scale and multi-resolution input for the model.
+The Decoders with cascaded skip connections module mainly achieves feature aggregation at skip connections and bridge layers in the cascaded network.
 The Aggregation module primarily reflects the aggregation of the decoder output layers and the application of our designed adaptive loss aggregation strategy.
 III-B Cascaded Backbones in MCFNet
-The cascaded backbone network consists of SEB (Sharp Extraction Backbone) and FCB (Flexible Connection Backbone).
-These two sub-backbone networks are responsible for feature extraction from images of different resolutions.
-Generally, high-resolution images can provide richer details and information, aiding the model in more accurately identifying and segmenting structures and regions within the images.
-However, this increases computational complexity and resource requirements.
-Low-resolution images are typically more robust to noise and interference but lose detailed information.
-Considering application requirements and computational resources, we choose to use FCB for processing $224\times 224$ images and SEB for assisting with the processing of $256\times 256$ images.
-In the encoder stage, the FCB sub-network uses traditional max pooling and convolution operations.
-This means that FCB has some limitations in feature extraction during the downsampling stage.
-Therefore, the design of SEB focuses on feature extraction during the downsampling stage while ensuring a relatively small number of parameters and computational load.
-Specifically, SEB enhances the recognition and extraction of detailed features during the downsampling stage for high-resolution images by adding SE attention mechanisms to the encoding layers.
-As the feature extraction process deepens in SEB, the images go through BatchNorm, MaxPool, and convolution operations via the SEB Stem and encoding layers.
-The prediction map of SEB is finally output from the highest layer of the SEB decoder and is subjected to a Hadamard product operation with the input image of SEB to obtain a new output image.
-This new output image is then resized to the $224\times 224$ input image required by FCB.
-The low-resolution images are input into FCB and processed through the FCB Stem and encoding layers.
-The four layers of feature maps are output through a skip connection mechanism and subsequently aggregated with those from SEB.
+The cascaded backbone network consists of SEB and FCB, responsible for extracting features from images of different resolutions.
+High-resolution images provide rich details but increase computational complexity.
+Low-resolution images are more robust to noise but lose details.
+We use FCB for $224\times 224$ images and SEB for $256\times 256$ images.
+FCB uses max pooling and convolution in the encoder, but the downsampling stage can further enhance feature capture.
+SEB focuses on improving feature extraction during downsampling with fewer parameters and computational load.
+Specifically, SEB enhances the extraction of detailed features during downsampling for high-resolution images by incorporating SE attention mechanisms in the encoding layers.
+The images pass through BatchNorm, MaxPool, and convolution operations in the SEB Stem and encoding layers.
+The final prediction map is produced from the SEB decoder’s highest layer, and a Hadamard product is applied with the input image to generate a new output image.
+This output is resized to the $224\times 224$ input required by FCB.
+Low-resolution images are processed in the FCB through its Stem and encoding layers, with feature maps output via skip-connections and subsequently aggregated with those from the SEB.
 III-C Cascaded Skip-connection Module in MCFNet
 As shown in Fig.
-2, this section mainly introduces the cascaded skip-connection mechanism.
-The efficient backbone model, FCB, is used to ensure the effectiveness of feature extraction, while a lightweight and simple SEB is employed to capture detailed information in high-resolution images, thus compensating for the limitations of FCB.
-Figure 2: (a) Illustration of the CSM (Cascaded Skip-connection Module). (b) Illustration of the LAT (Linear Attention Transformer Block) in the CSM.
-Specifically, when the encoding layers of FCB reach the lowest level after downsampling, they enter the CAB (Cross-Attention Bridge Layer).
-This module compensates for the loss of low-level features and then aggregates features with the output feature maps of the fourth encoding layer in SEB.
-For skip connections in FCB, the features first pass through the LAT (Linear Attention Transformer Block), enabling comprehensive and efficient feature extraction across both channel and spatial dimensions.
-As shown in Fig.
-2, the structure of LAT is based on our published work, FSCA-Net, specifically the PAT (Parallel Attention Transformer Block).
-We make new improvements to accelerate the convergence speed of the transformer-like structure and reduce the number of model parameters caused by its architecture.
-After the input tensor enters the LAT module, we first use a linear layer for dimensionality reduction to align the tensor dimensions with the channel numbers.
-The compressed dimensional feature maps are then processed sequentially through layers including LN, 2D-SCCA attention, and another LN layer.
-Finally, we obtain the output features by passing the data through a linear layer for dimensional restoration.
-Subsequently, the feature maps outputted by LAT are aggregated with the feature maps outputted by the skip connections of resized SEB.
-The aggregated feature maps are then fed into the decoding layer of FCB through skip connections, which directly propagate shallow features to deeper layers, preserving fine-grained information captured by shallow networks.
-In SEB, the output prediction map ($S_{output}$) from the highest decoding layer is fed into the SAM (Sigmoid Activation Map Module).
-This module mainly utilizes a 1x1 convolution to output a single-channel feature map, and then applies the $Sigmoid$ activation function to nonlinearly reweight features, enhancing the representation of key features while suppressing unimportant ones.
+4, this section introduces the cascaded skip-connection mechanism.
+The efficient FCB backbone ensures effective feature extraction, while the lightweight SEB captures detailed information in high-resolution images to complement FCB’s limitations.
+Specifically, when the encoding layers of FCB reach the lowest level after downsampling, they enter the Cross-Attention Bridge Layer proposed in our previous work, which compensates for the loss of low-level features and aggregates them with the output from SEB’s fourth encoding layer.
+FCB’s skip connections first pass through the Linear Attention Transformer Block (LAT), which enables efficient feature extraction across both channel and spatial dimensions.
+The LAT architecture is based on the parallel attention transformer module (PAT) from our previous work FSCA-Net, and has been improved to accelerate convergence and reduce the number of parameters.
+After the input tensor passes through the LAT module, it first undergoes dimensionality reduction via a linear layer to align the tensor dimensions with the channel numbers, reducing computational and storage demands.
+The compressed feature maps then pass through LN, 2D-SCCA attention, and another LN layer, ultimately being restored to their original dimensions through a linear layer.
+The feature maps from LAT are aggregated with those from resized SEB and passed into FCB’s decoding layer via skip connections, preserving shallow feature details.
+In SEB, the highest decoding layer’s output ($S_{output}$) is processed by the SAM module, which uses 1x1 convolution and the $Sigmoid$ activation function to enhance key features and suppress less important ones.
 The calculation formula for the SAM module is as follows:
@@
 $\begin{split}f(x,y)=&(1-\alpha)\cdot(1-\beta)\cdot pixel(i_{1},j_{1})\\ &+\alpha\cdot(1-\beta)\cdot pixel(i_{2},j_{2})\\ &+(1-\alpha)\cdot\beta\cdot pixel(i_{3},j_{3})\\ &+\alpha\cdot\beta\cdot pixel(i_{4},j_{4}),\end{split}$ (4)
-Figure 3: Illustration of the Aggregation module.
 where the target image coordinates are denoted as $(x,y)$, the original image coordinates are $(i,j)$, and the four nearest neighbor pixels in the original image are $(i_{1},j_{1})$, $(i_{2},j_{2})$, $(i_{3},j_{3})$, $(i_{4},j_{4})$.
 Additionally, $\alpha=x-i_{1}$, $\beta=y-j_{1}$, and $pixel(i,j)$ represent the pixel value at coordinates $(i,j)$ in the original image.
-Each decoding layer of SEB and FCB generates a prediction map, representing the predicted results for the input.
-These prediction maps come from different scales and levels, reflecting different levels of feature information and semantic understanding.
-We feed these prediction maps into the Aggregation module for the aggregation of multi-scale information, aiming to comprehensively utilize information from different scales and improve the accuracy and robustness of the final prediction results.
+Each decoding layer of SEB and FCB generates a prediction map at different scales, reflecting varying feature levels.
+These maps are input into the Aggregation module to combine multi-scale information, enhancing the accuracy and robustness of the final predictions.
+Figure 5: Illustration of the Aggregation module that merges hierarchical multi-scale features from FCB and SEB into a compact representation supervised by Adaptive-MFA.
 III-D Aggregation in MCFNet
 As shown in Fig.
-3, the module mainly summarizes the process of multi-scale feature aggregation in the decoding layer stage of two sub-networks, as well as the application of our newly designed adaptive loss aggregation strategy.
-Multi-scale feature aggregation helps the model utilize both global contextual information and local details, assisting the model in capturing finer image details and complex edge structures.
-This is particularly important for handling complex small structures in medical images, such as blood vessels and tumor boundaries.
-Moreover, merging features of different scales helps generate richer and more diverse feature representations, enabling the model to capture more comprehensive image information, effectively dealing with noise and deformation in images, and enhancing the segmentation capability for objects of different sizes.
-This allows the model to perform more stably across different types of medical images.
-Figure 4: Flowchart of the Adaptive-MFA Strategy.
-Specifically, the predicted maps output from the four layers of the encoding stage of SEB and B are pairwise fed into the corresponding convolutional blocks for channel-wise processing, obtaining feature maps with the desired specified number of channels.
+5, the module summarizes the process of multi-scale feature aggregation in the decoding stage and the application of our adaptive loss aggregation strategy.
+Multi-scale feature aggregation captures global context and fine details, improving segmentation of small, complex structures (e.g., blood vessels and tumors).
+By merging features from different scales, the model can handle noise, image deformations, and objects of varying sizes, ensuring more stable performance across diverse medical images.
+Specifically, the predicted maps output from the four layers of the encoding stage of SEB and FCB are pairwise fed into the corresponding convolutional blocks for channel-wise processing, obtaining feature maps with the desired specified number of channels.
 The calculation formula for this part is shown below:
 $O_{i}=Conv_{(C_{i},N,K)}(E_{i})(i=1,2,3,4),$ (5)
-where $E$ refers to the input feature map, $O$ refers to the output feature map, $C$ is the number of input channels, $N$ is the number of label classes, and $K$ is the convolution kernel size.
-Additionally, $C_{1}=64$, $C_{2}=64$, $C_{3}=128$, $C_{4}=256$.
-After the convolution block operations, the two sub-networks use bilinear interpolation resizing to match the dimensions of the output feature maps from FCB and SEB, followed by feature map aggregation.
-Subsequently, the four layers of different scale feature maps are fed into our designed adaptive loss aggregation strategy for prediction map data augmentation, thereby optimizing the training of the entire model.
-Through the adaptive search optimization accompanying the entire training process, the four layers of feature maps are aggregated to obtain the final output prediction map $Pred$ of the MCFNet model.
+where $E$ is the input feature map, $O$ is the output feature map, $C$ is the number of input channels, $N$ is the number of label classes, and $K$ is the convolution kernel size, with $C_{1}=64$, $C_{2}=64$, $C_{3}=128$, and $C_{4}=256$.
+After convolution, bilinear interpolation is used to resize feature maps from FCB and SEB for aggregation.
+The four scales of feature maps are then processed by an adaptive loss aggregation strategy for data augmentation, optimizing the training.
+The final output prediction map $Pred$ is obtained by aggregating these layers through adaptive search optimization during training.
 The aggregation of the four layers of feature maps can be expressed by the following mathematical formula:
 $Pred=u\times p_{1}+v\times p_{2}+w\times p_{3}+x\times p_{4},$ (6)
-where $p_{1}$, $p_{2}$, $p_{3}$, $p_{4}$ represent the feature maps from the four prediction heads, and $u$, $v$, $w$, $x$ are the corresponding weight coefficients for the four feature maps.
-In our experiments, we set all the weight coefficients to 1.
-In the following section, we will provide a detailed introduction to the adaptive loss aggregation strategy.
+where $p_{1}$, $p_{2}$, $p_{3}$, $p_{4}$ are the feature maps from the four prediction heads, with $u$, $v$, $w$, $x$ as their corresponding weight coefficients, all set to 1 in our experiments.
+Our model produces accurate segmentation maps.
+Within the aggregation module, combining multi-scale features refines boundaries and improves localization, thereby enhancing overall prediction performance and clinical utility.
+Figure 6: Flowchart of the Adaptive-MFA Strategy, which is executed concurrently with the training process.
 III-E Adaptive-MFA: Adaptive Multi-scale Feature-Mixing Loss Aggregation
-To better optimize model training, we have designed an adaptive loss aggregation strategy.
-The main idea is to create new prediction maps by combining available prediction maps.
-By integrating various prediction maps, the model can better handle errors in different situations.
-Adaptively selecting and weighting different prediction maps can optimize the final prediction results based on performance in different scenarios.
-This allows the model to better capture complex patterns in the data, thereby improving prediction accuracy.
-The entire strategy flowchart is illustrated in the Fig.
-4.
-Specifically, we randomly combine the output prediction maps $P_{1}$, $P_{2}$, $P_{3}$ and $P_{4}$ from the four layers of the decoder stage, resulting in a total of 15 non-empty subset combinations.
-Subsequently, we divide the 15 combinations into four sets $S_{1}$, $S_{2}$, $S_{3}$ and $S_{4}$, and assign corresponding weights $W_{1}$, $W_{2}$, $W_{3}$ and $W_{4}$ to these sets.
-These weight coefficients are adaptively adjusted during model training.
-Regarding the setting of weight coefficients, since $W_{1}+W_{2}+W_{3}+W_{4}=1$ imposes a competitive relationship among the four weight coefficients, it restricts the optimization of model training as it cannot account for all prediction map combinations.
-In other words, some prediction maps might contribute minimally to the optimization of the training loss but are still crucial for the model training process.
-Therefore, we discard the $W_{1}+W_{2}+W_{3}+W_{4}=1$ restriction and only stipulate that the initial values satisfy $W_{1}=W_{2}=W_{3}=W_{4}$.
+To optimize model training, we design an adaptive loss aggregation strategy, primarily by combining different prediction maps to create new ones.
+The adaptive selection and weighting of prediction maps help optimize the final results for better prediction accuracy in varying scenarios.
+The strategy flowchart in Fig.
+6 shows how we combine output prediction maps $P_{1}$, $P_{2}$, $P_{3}$, and $P_{4}$ from the decoder’s four layers, forming 15 non-empty subset combinations.
+These are divided into four sets $S_{1}$, $S_{2}$, $S_{3}$, with adaptive weights $W_{1}$, $W_{2}$, $W_{3}$, and $W_{4}$ adjusted during training.
+We remove the constraint $W_{1}+W_{2}+W_{3}+W_{4}=1$ because it hinders prediction maps with minimal contribution from playing an important role in model training.
+Initially, all weights are set equal: $W_{1}=W_{2}=W_{3}=W_{4}$.
 The specific contents of the four sets are as follows:
@@
 $\begin{split}S_{2}=&\Big\{\{p_{1},p_{2}\},\{p_{1},p_{3}\},\{p_{1},p_{4}\},\\ &\{p_{2},p_{3}\},\{p_{2},p_{4}\},\{p_{3},p_{4}\}\Big\},\end{split}$ (8)
-$\begin{split}S_{3}=&\Big{\{}\{p_{1},p_{2},p_{3}\},\{p_{1},p_{2},p_{4}\},\\ &\{p_{1},p_{3},p_{4}\},\{p_{2},p_{3},p_{4}\}\Big{\}},\\ \end{split}$ (9)
+[…]
 $S_{4}=\Big\{\{p_{1},p_{2},p_{3},p_{4}\}\Big\}.$ (10)
@@
 IV Experiments
-This section first introduces all 10 datasets used in the experiment, along with the quantitative and qualitative analyses of the experiment.
-Subsequently, we provide a detailed description of the parameter configurations used in the experiment.
-Finally, we outline the ablation study conducted for our proposed method.
+This section introduces the datasets and implementation settings, then reports detailed quantitative and qualitative results, followed by thorough ablation studies and discussion of the proposed method.
 IV-A Datasets
-To validate the efficient segmentation of whole-body risk organs and lesions by MCFNet, we conduct experiments on ten datasets. The detailed descriptions of each dataset are as follows:
+To assess whole-body OARs and lesion segmentation, we evaluate MCFNet on ten datasets spanning multiple anatomical regions and imaging protocols.
+For each dataset, we report the source, data volume, and annotation scope.
 IV-A1 HaN (2015)
-This is a public dataset for risk organs in the head and neck region, containing CT and MRI images of patients’ head and neck.
-These patients underwent CT scans and T1-weighted MR imaging to support image-guided radiotherapy planning.
-The segmentation labels provided by the dataset are stored separately; we preprocessed the segmentation labels of six regions, merging them into one segmentation label file and classifying them in the order of brainstem, mandible, left parotid gland, right parotid gland, left submandibular gland, and right submandibular gland.
-After preprocessing the dataset, we divide the image data of 17 tumor patients into 14 cases for the training set and 3 cases for the test set, with the training set containing 1924 slices.
+This public dataset contains CT and MRI images of patients’ head and neck, used for image-guided radiotherapy planning.
+The dataset provides segmentation labels for six regions: brainstem, mandible, left and right parotid gland, left and right submandibular gland.
+After preprocessing, we merged these labels into a single file.
+The dataset consists of images from 17 tumor patients, with 14 cases used for training (1924 slices) and 3 cases for testing.
 IV-A2 SegTHOR (2019)
-This dataset is a CT dataset specifically for the segmentation of thoracic risk organs (OARs) and was included in the ISBI 2019 challenge.
-In this dataset, the risk organs for the segmentation task include the heart, trachea, aorta, and esophagus, which differ in spatial and appearance characteristics.
-After preprocessing the dataset, we divide the image data of 40 cases into 32 cases for the training set and 8 cases for the test set, with the training set containing 6035 slices.
+This CT dataset, part of the ISBI 2019 challenge, is designed for thoracic risk organ segmentation, including the heart, trachea, aorta, and esophagus.
+These organs differ in spatial and appearance characteristics.
+After preprocessing, we divided the dataset of 40 cases into 32 for training (6035 slices) and 8 for testing.
 IV-A3 CHAOS (2019)
-As a multi-modal abdominal medical image segmentation dataset, it provides paired multi-modal CT and MRI data along with corresponding annotation information.
-MRI imaging data is obtained from two different sequences (T1-DUAL and T2-SPIR).
-Since the CT images only provide liver annotations, and the MRI images provide annotations for four organs (liver, spleen, left kidney, and right kidney), we select MRI imaging of 20 patients with two modalities.
-After preprocessing the dataset, we divide the image data of 40 cases into 32 cases for the training set and 8 cases for the test set, with the training set containing 1026 slices.
+This multimodal abdominal segmentation dataset comprises paired CT–MRI volumes with corresponding annotations.
+MRI data includes T1-DUAL and T2-SPIR sequences, with annotations for liver, spleen, left kidney, and right kidney.
+After preprocessing, we split the data of 40 cases into 32 for training (1026 slices) and 8 for testing.
 IV-A4 LiTS (2017)
-This dataset focuses on the segmentation of the liver and its tumors in CT images.
-The data were collected from seven different medical centers and were included in the ISBI 2017, MICCAI 2017, and MICCAI 2018 segmentation challenges.
-From the original dataset of 131 patients, we randomly select data from 40 patients.
-After preprocessing the images, we obtain 32 cases for the training set and 8 cases for the test set, with the training set containing 4599 slices.
+This dataset focuses on liver and tumor segmentation in CT images.
+It was collected from seven medical centers and included in the ISBI 2017, MICCAI 2017, and MICCAI 2018 challenges.
+From 131 patients, we randomly select data from 40, with 32 cases for training (4599 slices) and 8 cases for testing.
 IV-A5 KiTS (2019)
-This dataset is a CT dataset for the segmentation of kidneys and their tumors.
-It was included in the MICCAI 2019 challenge, which aimed to accelerate the development of reliable semantic segmentation methods for kidneys and kidney tumors.
-All data were sourced from arterial phase abdominal CT scans of 300 kidney cancer patients who underwent partial or radical nephrectomy, with ground truth labels for kidneys and tumors provided.
-From the 210 publicly available cases, we randomly select data from 80 patients.
-After preprocessing the images, we obtain 64 cases for the training set and 16 cases for the test set, with the training set containing 4754 slices.
+This dataset focuses on kidney and tumor segmentation in CT images and was part of the MICCAI 2019 challenge.
+It includes data from 300 kidney cancer patients who underwent nephrectomy.
+From 210 available cases, we randomly select 80, with 64 cases for training (4754 slices) and 16 cases for testing.
 IV-A6 KiTS (2023)
-This dataset focuses on the segmentation of kidneys, tumors, and cysts, and is an extension of the KiTS (2019) dataset.
-It includes additional segmentation labels for cysts and increases the total data volume from 210 cases to 489 cases.
-This dataset was included in the MICCAI 2023 challenge.
-From the 489 patients’ image data, we randomly select data from 80 patients for model validation.
-After preprocessing the data, we divide it into 64 cases for the training set and 16 cases for the test set, with the training set containing 4968 slices.
+This dataset extends the KiTS (2019) dataset to include cyst segmentation, increasing the total cases from 210 to 489.
+It was part of the MICCAI 2023 challenge.
+From 489 cases, we randomly select 80, with 64 cases for training (4968 slices) and 16 cases for testing.
 IV-A7 LUNG (2018)
-This dataset is a subtask of the MSD Medical Image Segmentation Decathlon Challenge, aiming to segment lung tumors from CT images, highlighting the challenge of segmenting small targets within a large context.
-The official dataset includes thin-slice CT scans from 63 non-small cell lung cancer patients.
-After preprocessing the images, we divide the dataset into 50 cases for the training set and 13 cases for the test set, with the training set containing 1264 slices.
-IV-A8 CPCGEA (2023)
-This dataset includes MRI data of prostate cancer patients, comprising two modalities: DWI and T2WI sequences.
-The images suffer from poor clarity and additional noise due to environmental conditions or image sensor issues.
-Errors during image acquisition can also lead to artifacts such as streaks and shadows.
-The dataset consists of 172 cases with manually annotated prostate cancer regions.
-In our experiments, we randomly divide the data into 139 cases for the training set and 33 cases for the test set, with the training set containing 832 slices.
-IV-A9 KNEE (2023)
-This dataset consists of MRI images of patients’ leg bones, focusing on the segmentation of the femur.
-It includes MRI image data from 59 patients.
-After preprocessing the images, we obtain 47 cases for the training set and 12 cases for the test set, with the training set containing 8171 slices.
+This dataset is part of the MSD Medical Image Segmentation Decathlon Challenge, focusing on lung tumor segmentation from CT scans.
+It includes thin-slice CT scans from 63 non-small cell lung cancer patients.
+After preprocessing, we divide it into 50 cases for training (1264 slices) and 13 cases for testing.
 ((a))
@@
 ((d))
-Figure 5: Training loss trends of the proposed method and seven comparative methods across four datasets.
+Figure 7: Training loss curves of the proposed method and seven comparison methods across four datasets. The stable downward trend of the loss demonstrates the orderly progression of model training.
+IV-A8 CPCGEA (2023)
+This dataset contains MRI data of prostate cancer patients, including DWI and T2WI sequences.
+The images are affected by poor clarity and noise, with artifacts like streaks and shadows due to image acquisition issues.
+It includes 172 cases with annotated prostate cancer regions.
+We randomly split the data into 139 training cases (832 slices) and 33 test cases.
+IV-A9 KNEE (2024)
+This dataset consists of MRI images of patients’ leg bones, focusing on femur segmentation.
+It includes data from 59 patients.
+After preprocessing, we use 47 cases for training (8171 slices) and 12 cases for testing.
 IV-A10 AutoPET (2023)
-AutoPET is a large-scale PET/CT dataset focusing on whole-body tumor segmentation, included in the MICCAI 2022 and MICCAI 2023 challenges.
-The official dataset comprises 1014 paired PET-CT images from 900 patients.
-Each data instance includes a three-dimensional whole-body FDG-PET image, a three-dimensional whole-body CT image, and a manually annotated three-dimensional tumor mask.
-Notably, all CT and PET images were acquired in the same scanning session using the same PET/CT scanner, ensuring high anatomical consistency.
-The dataset primarily covers patients with histologically confirmed malignant melanoma, lymphoma, and lung cancer, as well as negative control patients who underwent FDG-PET/CT scans at the University Hospital of Tübingen and LMU University Hospital in Munich, Germany.
-In our experiments, we select the PET images and annotations.
-We randomly divide the data from 80 patients into 64 cases for the training set and 16 cases for the test set, with the training set containing 2558 slices.
+AutoPET is a large-scale PET/CT dataset for whole-body tumor segmentation, included in the MICCAI 2022 and MICCAI 2023 challenges.
+It contains 1014 paired PET-CT images from 900 patients, focusing on malignant melanoma, lymphoma, and lung cancer.
+The dataset includes three-dimensional FDG-PET and CT images, with manually annotated tumor masks.
+We select PET images and annotations from 80 patients, dividing the data into 64 training cases and 16 test cases, with 2558 slices in the training set.
 IV-B Evaluation Metrics
-In all experiments, we comprehensively evaluate the performance of the segmentation models by applying appropriate metrics based on the dataset type.
-For segmentation tasks on multi-class datasets, we use two key evaluation metrics: the Dice Similarity Coefficient (DSC) and the 95% Hausdorff Distance (HD95).
-DSC measures the overlap between the predicted segmentation and the ground truth, with higher values indicating better segmentation performance.
-HD95 assesses the boundary deviations in the segmentation results, where smaller values reflect higher accuracy in boundary prediction.
-For segmentation tasks on single-class datasets, we use four evaluation metrics: DSC, HD95, Recall, and Precision.
-In addition to DSC and HD95, Recall measures the model’s ability to detect all instances of the target class, showing how well the model captures the entire target region.
-Precision measures the proportion of correctly identified target regions among all predicted positives, evaluating the model’s ability to avoid false positives.
-By combining these four metrics, we can more comprehensively and accurately assess the performance of single-class segmentation models.
-Figure 6: Visual examples of segmentation on the HaN, SegTHOR, CHAOS, and LiTS datasets.
+We evaluate segmentation models using appropriate metrics based on the dataset type.
+For multi-class segmentation, we evaluate using the Dice similarity coefficient (DSC) and the 95th-percentile Hausdorff distance (HD95).
+DSC measures overlap between the prediction and ground truth (higher is better), while HD95 quantifies boundary error at the 95th percentile (lower is better).
+For single-class segmentation, we report DSC, HD95, Recall (Sensitivity), and Precision (Positive Predictive Value).
+Recall evaluates the model’s ability to capture all target voxels (penalizing false negatives), and Precision measures the proportion of predicted positives that are correct (penalizing false positives).
+Together, these metrics provide a balanced evaluation of volumetric agreement and boundary accuracy.
+Figure 8: Visual examples of segmentation on the HaN, SegTHOR, CHAOS, and LiTS datasets. In the image, different organs or lesions are highlighted using distinct color annotations.
 IV-C Implementation Detail
@@
 $\begin{split}L(Y,P)=&1-\sum_{i=1}^{I}(\lambda\frac{2*\sum_{n=1}^{N}{Y}_{n,i}\cdot{P}_{n,i}}{\sum_{n=1}^{N}{Y}_{n,i}^{2}+\sum_{n=1}^{N}{P}_{n,i}^{2}}\\ &+\sum_{n=1}^{N}{Y}_{n,i}log{P}_{n,i}),\end{split}$ (12)
-$I$ is the number of classes, $N$is the total number of voxel. ${Y}_{n,i}$ and ${P}_{n,i}$ are the ground truth and output probability for the $i$-th class at voxel $n$, respectively.
-IV-D Comparsion With Other Methods
-We employ seven state-of-the-art methods to validate the performance of MCFNet: U-Net, TransUNet, Swin-Unet, UCTransNet, TransCASCADE, Cascaded MERIT, and FSCA-Net.
-We validate the segmentation performance of our proposed method against seven comparison models across all ten datasets.
-In Fig.
-5, the training loss trends of all methods are illustrated across the four datasets.
-Fig.
-5 shows visual examples of segmentation on all datasets and we highlight incorrectly segmented areas with red boxes.
-The specific qualitative and quantitative analysis results are as follows:
-IV-D1 Experiments on HaN
-Our proposed method and seven comparison models segment six regions in this dataset.
-As shown in the Table I, the Cascaded MERIT model achieves excellent segmentation performance, with an average DSC of 74.71% and an average HD95 of 2.532mm.
-Our designed MCFNet improves the average DSC by 2.81% and reduces the average HD95 by 0.079mm compared to Cascaded MERIT.
-The experimental results in Table I show that our model significantly improves the segmentation of the left submandibular gland.
-As shown in the first row of Fig.
-6, we observe that our proposed method does not exhibit under-segmentation or over-segmentation compared to the comparison models.
-For example, UNet incorrectly identifies the mandibular region, while Swin-Unet fails to fully recognize the brainstem region, resulting in under-segmentation.
-Additionally, UCTransNet fails to accurately distinguish between the left and right parotid glands.
+II is the number of classes, NNis the total number of voxel. ${Y}_{n,i}$ and ${P}_{n,i}$ are the ground truth and output probability for the $i$-th class at voxel $n$, respectively.
 TABLE I: COMPARISONS WITH STATE-OF-THE-ART MODELS ON THE HAN DATASET.
@@
 Ours 77.52 2.453 79.52 93.47 77.59 78.61 68.02 67.92
+IV-D Comparsion With Other Methods
+We validate the performance of MCFNet across ten datasets using seven state-of-the-art methods: U-Net, TransUNet, Swin-Unet, UCTransNet, TransCASCADE, Cascaded MERIT, and FSCA-Net.
+Fig.
+7 shows the training loss trends of all methods across four datasets, while Fig.
+8, 9, 10, and 11 display visual segmentation examples for each dataset, with errors highlighted in red boxes.
+The specific qualitative and quantitative analysis results are as follows:
+IV-D1 Experiments on HaN
+Our proposed method and seven comparison models segment six regions in this dataset.
+As shown in the Table I, the Cascaded MERIT model achieves excellent segmentation performance, with an average DSC of 74.71% and an average HD95 of 2.532mm.
+Our designed MCFNet improves the average DSC by 2.81% and reduces the average HD95 by 0.079mm compared to Cascaded MERIT.
+The results in Table I indicate that our model achieves a substantial performance gain on the left submandibular gland.
+In the first row of Fig.
+8, we see that our method avoids both under-segmentation and over-segmentation.
+However, UNet misidentifies the mandibular region, Swin-Unet under-segments the brainstem, and UCTransNet struggles to distinguish between the left and right parotid glands.
+IV-D2 Experiments on SegTHOR
+Our proposed method and seven comparison models segment the trachea, heart, aorta, and esophagus.
+As shown in Table II, state-of-the-art methods like TransUNet, UCTransNet, and FSCA-Net achieve over 80% DSC and reduce the average HD95 to below 5mm.
+Our method achieves an average DSC of 83.27% and HD95 of 4.187mm, outperforming others in segmenting these organs.
+As shown in the second row of Fig.
+8, our designed model accurately identifies and segments the risk organs.
+However, UNet, TransUNet, UCTransNet, Cascaded MERIT, and FSCA-Net all mistakenly identify the background as the aorta.
+Additionally, Swin-Unet incorrectly identifies the esophagus region as the heart.
+IV-D3 Experiments on CHAOS
+The comparison results are shown in Table III.
+We can see that the cascade networks TransCASCADE [40] and Cascaded MERIT achieve relatively good segmentation results for the four abdominal organs, with average DSCs of 90.26% and 90.69%, and average HD95s of 3.695mm and 3.914mm, respectively.
+The traditional TransUNet achieves an average DSC of 91.19%, making it the best-performing model among the comparisons.
+Achieving 92.80% average DSC and 3.676 mm average HD95, the proposed method surpasses all competing models.
+As shown in the third row of Fig.
+8, most comparison models under-segment the liver, and Cascaded MERIT also struggles with identifying the spleen and right kidney.
+Our model achieves generally accurate segmentation, but the spleen’s edges could still be improved.
+IV-D4 Experiments on LiTS
+All results for this dataset are shown in Table IV.
+We can see that TransCASCADE is the best-performing advanced segmentation method, achieving an average DSC of 72.56%.
+For liver tumor segmentation, TransUNet achieves a DSC of 51.93%, which is currently the best.
+Our designed model achieves better segmentation results for tumors, with an average DSC of 73.79% and a DSC of 53.35% for tumor regions.
 TABLE II: COMPARISONS WITH STATE-OF-THE-ART MODELS ON THE SEGTHOR DATASET.
@@
 Ours 92.80 3.676 94.87 93.45 92.05 90.83
-IV-D2 Experiments on SegTHOR
-Our proposed method and seven comparison models segment the trachea, heart, aorta, and esophagus in this dataset.
-According to the experimental results in Table II, current state-of-the-art methods such as TransUNet, UCTransNet, and FSCA-Net achieve over 80% in average DSC and reduce the average HD95 to below 5mm.
-Our proposed method achieves an average DSC of 83.27% and an average HD95 of 4.187mm, with MCFNet obtaining the best performance in segmenting the trachea, heart, aorta, and esophagus.
-As shown in the second row of Fig.
-6, our designed model accurately identifies and segments the risk organs.
-However, UNet, TransUNet, UCTransNet, Cascaded MERIT, and FSCA-Net all mistakenly identify the background as the aorta.
-Additionally, Swin-Unet incorrectly identifies the esophagus region as the heart.
-IV-D3 Experiments on CHAOS
-The comparison results are shown in Table III.
-We can see that the cascade networks TransCASCADE [33] and Cascaded MERIT achieve relatively good segmentation results for the four abdominal organs, with average DSCs of 90.26% and 90.69%, and average HD95s of 3.695mm and 3.914mm, respectively.
-The traditional TransUNet achieves an average DSC of 91.19%, making it the best-performing model among the comparisons.
-Our designed model achieves an average DSC of 92.80% and reduces the average HD95 to 3.676mm, significantly outperforming all comparison models.
-As shown in the third row of Fig.
-6, most comparison models exhibit under-segmentation for the liver region, and Cascaded MERIT also shows ambiguous identification for the spleen and right kidney.
-Our designed model generally achieves accurate segmentation, but there is still room for improvement in the segmentation of the spleen’s edges.
 TABLE IV: COMPARISONS WITH STATE-OF-THE-ART MODELS ON THE LITS DATASET.
@@
 Ours 73.79 23.326 94.22 53.35
+As shown in the fourth row of Fig.
+8, our model provides accurate segmentation for the liver and tumors.
+However, fine delineation of tumor boundaries remains challenging.
+Models such as UNet often struggle to detect small lesions, and even Cascaded MERIT and FSCA-Net exhibit suboptimal performance in capturing hepatic boundary details.
+IV-D5 Experiments on KiTS (2019)
+The dataset targets segmentation of the kidneys and renal tumors.
+As shown in Table V, our designed model achieves an average DSC that is 2.21% higher than the state-of-the-art model Cascaded MERIT and an average HD95 that is 2.448mm lower.
+Additionally, MCFNet improves tumor segmentation DSC by 4.49% over TransCASCADE, and achieves a further 0.21% DSC gain on kidney segmentation compared with Cascaded MERIT.
+As shown in the first row of Fig.
+9, we show the segmentation effects on an image containing both the kidney and tumor.
+Swin-Unet has scattered segmentation in the tumor region, UCTransNet mistakenly identifies some background areas as tumors, and FSCA-Net fails to recognize the tumor regions.
 TABLE V: COMPARISONS WITH STATE-OF-THE-ART MODELS ON THE KITS (2019) DATASET.
@@
 Ours 70.71 20.522 90.50 50.92
-IV-D4 Experiments on LiTS
-All results for this dataset are shown in Table IV.
-We can see that TransCASCADE is the best-performing advanced segmentation method, achieving an average DSC of 72.56%.
-For liver tumor segmentation, TransUNet achieves a DSC of 51.93%, which is currently the best.
-Our designed model achieves better segmentation results for tumors, with an average DSC of 73.79% and a DSC of 53.35% for tumor regions.
-As shown in the fourth row of Fig.
-6, our designed model, like many comparison models, achieves relatively accurate segmentation for the liver and liver tumors.
-However, there is a gap in the detailed segmentation of the tumor region edges.
-Additionally, models like UNet, due to the small size of the segmentation target, may fail to recognize the tumor region.
-Figure 7: Visual examples of segmentation on the KiTS (2019), KiTS (2023), LUNG, CPCGEA and KNEE datasets.
 TABLE VI: COMPARISONS WITH STATE-OF-THE-ART MODELS ON THE KITS (2023) DATASET.
@@
 Ours 64.07 22.770 91.74 45.18 55.31
-IV-D5 Experiments on KiTS (2019)
-This dataset focuses on the segmentation of kidneys and kidney tumors.
-As shown in Table V, our designed model achieves an average DSC that is 2.21% higher than the state-of-the-art model Cascaded MERIT and an average HD95 that is 2.448mm lower.
-Additionally, MCFNet improves the tumor segmentation DSC by 4.49% compared to TransCASCADE.
-As shown in the fifth row of Fig.
-7, we demonstrate the segmentation effects of the models on an image containing both the kidney and tumor.
-Swin-Unet exhibits scattered segmentation in the tumor region, with dispersed pixels.
-UCTransNet mistakenly identifies some background areas as tumors.
-Even more problematic, FSCA-Net fails to recognize the tumor region altogether.
 TABLE VII: COMPARISONS WITH STATE-OF-THE-ART MODELS ON THE LUNG DATASET.
@@
 Ours 74.49 23.505 71.06 83.52
-Figure 8: Visual examples of segmentation on the AutoPET dataset.
+Figure 9: Visual examples of segmentation on the KiTS (2019), KiTS (2023), LUNG, CPCGEA and KNEE datasets. In the image, different organs or lesions are highlighted using distinct color annotations.
 IV-D6 Experiments on KiTS (2023)
-This dataset builds upon the segmentation labels of kidneys and kidney tumors from KiTS (2019) by adding segmentation for cysts.
-As shown in Table VI, traditional networks do not perform well in the segmentation of tumors and cysts.
-For example, the best-performing TransUNet achieves an average DSC of only 42.31% for tumors and 42.68% for cysts.
-In contrast, our designed model improves the DSC for cyst segmentation to 55.31%, and MCFNet achieves the best performance in segmentation for kidneys, kidney tumors, and cysts.
+As shown in Table VI, traditional networks perform poorly in tumor and cyst segmentation, with the best TransUNet achieving a DSC of only 42.31% for tumors and 42.68% for cysts.
+In contrast, our model improves the DSC for cyst segmentation to 55.31%, and MCFNet excels in segmenting kidneys, tumors, and cysts.
+As shown in the second row of Fig.
+9, cyst segmentation is more challenging due to small target areas.
+Models like TransUNet, and Swin-Unet struggle with cyst segmentation.
+While UCTransNet correctly identifies cyst regions, it has difficulty maintaining accurate tumor segmentation simultaneously.
 TABLE VIII: COMPARISONS WITH STATE-OF-THE-ART MODELS ON THE CPCGEA DATASET.
@@
 Ours 99.14 1.000 98.99 99.29
+IV-D7 Experiments on LUNG
+We evaluate segmentation performance using four metrics, as summarized in Table VII.
+Among the baselines, FSCA-Net achieves the best DSC and Recall, while Cascaded MERIT excels in Precision at 82.23%.
+Our model outperforms all models in DSC, HD95, Recall, and Precision.
+Specifically, MCFNet improves DSC by 2.39% and Recall by 2.99% compared to FSCA-Net.
+As shown in the third row of Fig.
+9, most models correctly identify tumor regions, except Swin-Unet.
+UNet includes background areas as tumors, while Swin-Unet misidentifies both tumor and vascular regions as tumors.
+Figure 10: Visual examples of segmentation on the AutoPET dataset. The tumor lesion region is marked in red for clarity.
+Figure 11: The visual results of our method on the AutoPET dataset. The tumor lesion region is marked in red for clarity.
 TABLE X: COMPARISONS WITH STATE-OF-THE-ART MODELS ON THE AUTOPET DATASET.
@@
 Ours 63.87 22.251 65.35 73.91
-As shown in the sixth row of Fig.
-7, the addition of cyst segmentation labels increases the difficulty of segmenting small target areas.
-Models like UNet, TransUNet, and Swin-Unet demonstrate shortcomings in segmenting cysts effectively.
-Additionally, while UCTransNet accurately identifies cyst regions, it struggles to maintain accurate recognition of tumor regions simultaneously.
-IV-D7 Experiments on LUNG
-In this dataset, we select four evaluation metrics to validate the model’s segmentation performance, as shown in Table VII.
-It is observed that the advanced model FSCA-Net achieves the best results in terms of DSC and Recall, while Cascaded MERIT achieves the best Precision among existing models, reaching 82.23%.
-Our designed model achieves the best performance across DSC, HD95, Recall, and Precision metrics.
-Specifically, MCFNet improves DSC by a significant 2.39% and Recall by 2.99% compared to FSCA-Net.
-As shown in the seventh row of Fig.
-7, for single-label segmentation targets, most models correctly identify tumor regions except for Swin-Unet.
-UNet identifies tumor regions but also mistakenly includes background areas as tumors.
-Swin-Unet not only fails to accurately identify tumor regions but also misidentifies vascular regions in the background as tumors.
 IV-D8 Experiments on CPCGEA
-As shown in Table VIII, we compare seven comparison models for the segmentation performance of prostate cancer.
-Since we are dealing with single-label image segmentation, we use four evaluation metrics to comprehensively assess model performance.
-Our designed model improves upon the current best model by 1.78% in DSC, reduces HD95 by 0.582mm, increases Recall by 1.01%, and improves Precision by 0.29%.
-As shown in the eighth row of Fig.
-7, we continue the work initiated by our previously proposed model, FSCA-Net, and conducted further exploration on CPCGEA.
-We add experiments with TransCASCADE and Cascaded MERIT, revealing that they exhibit unclear boundary segmentation for prostate cancer regions and struggle to delineate the edges of tumor areas effectively.
-TABLE XI: ABLATION EXPERIMENTS ON FIVE DATASETS.
+As shown in Table VIII, we evaluate seven models for prostate cancer segmentation using four metrics.
+The proposed method attains the best overall performance, exceeding the previous state of the art by 1.78% in DSC, reducing HD95 by 0.582 mm, and yielding gains of 1.01% in Recall and 0.29% in Precision.
+As shown in the fourth row of Fig.
+9, the comparative models fail to clearly delineate the boundaries of the tumor region.
+TransCASCADE and Cascaded MERIT are prone to mistakenly identifying background areas as lesions.
+IV-D9 Experiments on KNEE
+As shown in Table IX, all models achieve an average DSC for femur segmentation exceeding 97.5%, with MCFNet achieving the highest at 99.14%.
+Furthermore, HD95, Recall, and Precision also achieve optimal performance, although the improvement in segmentation performance compared to advanced models is relatively limited.
+As depicted in the fifth row of Fig.
+9, we can see that all models accurately segment the femur region.
+However, Swin-Unet and UCTransNet exhibit minor segmentation errors in very small areas, although the overall differences are minimal.
+IV-D10 Experiments on AutoPET
+For whole-body tumor segmentation, our designed model and seven comparison models are evaluated using four metrics, as shown in the Table X.
+TransUNet achieves the best DSC of 61.95%, the best HD95 of 22.329mm, and the best Precision of 72.98%.
+Additionally, FSCA-Net achieves a Recall of 62.07%.
+However, our designed model MCFNet improves by 1.92% in DSC, reduces HD95 by 0.078mm, increases Recall by 3.28%, and improves Precision by 0.93%.
+Fig.
+10 and Fig.
+11 present four segmentation cases from the AutoPET dataset, with erroneous areas marked in red boxes.
+In the whole-body tumor segmentation images, there are many background areas that appear white and have smooth edges but are not tumors, leading the segmentation models to mistakenly identify tumor regions.
+Most models confuse background with tumor regions, with UNet and TransUNet showing over- and under-segmentation.
+Swin-Unet, UCTransNet, and FSCA-Net also exhibit errors, while TransCASCADE and Cascaded MERIT show under-segmentation.
+MCFNet performs well but can still improve in segmenting detailed tumor boundaries.
+TABLE XI: ABLATION EXPERIMENTS ON SEGTHOR, CHAOS, LITS, KITS2019, AND CPCGEA DATASETS.
 Method SegTHOR CHAOS LiTS KiTS19 CPCGEA
@@
 SEB+FCB+Adaptive-MFA 83.27 4.187 92.80 3.676 73.79 23.326 70.71 20.522 67.28 6.064 71.48 69.43
-Figure 9: The experiments shown in the figure are conducted on the CPCGEA dataset. (a) Visualization of model performance and parameter size. (b) Visualization of model performance and FLOPs.
-IV-D9 Experiments on KNEE
-As shown in Table IX, all models achieve an average DSC for femur segmentation exceeding 97.5%, with our designed model achieving the highest at 99.14%.
-Additionally, HD95, Recall, and Precision also achieve optimal performance, although the improvement in segmentation performance compared to advanced models is relatively limited.
-As depicted in the ninth row of Fig.
-7, we can see that all models accurately segment the femur region.
-However, Swin-Unet and UCTransNet exhibit minor segmentation errors in very small areas, although the overall differences are minimal.
-IV-D10 Experiments on AutoPET
-For whole-body tumor segmentation, our designed model and seven comparison models are evaluated using four metrics, as shown in the Table X.
-Among the current advanced methods, TransUNet achieves the best DSC of 61.95%, the best HD95 of 22.329mm, and the best Precision of 72.98%.
-Additionally, FSCA-Net achieves a Recall of 62.07%.
-However, our designed model MCFNet improves by 1.92% in DSC, reduces HD95 by 0.078mm, increases Recall by 3.28%, and improves Precision by 0.93%.
-As shown in Fig.
-8, four segmentation cases from the AutoPET dataset are presented.
-The erroneous segmentation areas are all marked with red boxes.
-In the whole-body tumor segmentation images, there are many background areas that appear white and have smooth edges but are not tumors, leading the segmentation models to mistakenly identify tumor regions.
-From a comprehensive observation of the four cases, it can be seen that most models exhibit confusion between the background and tumor regions.
-Specifically, UNet and TransUNet have issues with over-segmentation and under-segmentation.
-Swin-Unet, UCTransNet, and FSCA-Net all display errors in segmentation, while TransCASCADE and Cascaded MERIT show cases of under-segmentation.
-Although the MCFNet model has performed relatively well in tumor segmentation, there is still room for improvement in segmenting the detailed boundaries of tumor regions.
+TABLE XII: COMPLEXITY COMPARISON BETWEEN MLP AND LINEAR LAYERS IN THE LAT MODULE.
+[…]
+(G) (M)
+LAT w/ MLP 61.2 57.7
+LAT w/ Linear 59.3 47.0
 IV-E Ablation Study
@@
 IV-F Discussion
+IV-F1 Discussion on the Complexity Optimization of the LAT Module
+As shown in Table XII, replacing the MLP layer with a linear layer in the LAT module effectively reduces model complexity.
+Specifically, the number of parameters decreases from 57.7M to 47.0M, and the computational cost (FLOPs) is reduced from 61.2G to 59.3G.
+The structural simplification is achieved by removing redundant non-linear transformations present in the MLP.
+Consequently, this optimization yields a more efficient and practical LAT module, making it particularly suitable for resource-constrained application scenarios.
+TABLE XIII: ACCURACY (DICE SCORE) VS. MODEL COMPLEXITY (PARAMETERS AND COMPUTATIONAL COMPLEXITY) COMPARISON ON THE CPCGEA DATASET.
+[…]
+(%, mean) (G) (M)
+U-Net 64.1 50.3 14.8
+TransUNet 63.2 56.7 105.0
+SwinUnet 60.9 67.3 82.3
+UCTransNet 60.3 63.2 65.6
+TransCASCADE 65.4 22.1 123.5
+Cascaded MERIT 65.5 33.3 147.9
+FSCA-Net 65.2 32.7 44.2
+Ours 67.3 59.3 47.0
+IV-F2 Discussion on the Model Complexity
 MCFNet integrates low parameter count, low computational cost, and high segmentation performance.
 Our designed model effectively balances performance with parameter and computational cost, demonstrating significant advantages over existing state-of-the-art methods.
-To illustrate this, we conduct experiments on the CPCGEA dataset and compare our model’s results with several benchmark models.
-These experimental results are visualized in figures, clearly showing the performance of the MCFNet model.
-As shown in Fig.
-9, our model not only exhibits excellent segmentation performance but also achieves or surpasses the performance of existing methods while maintaining a smaller parameter size and lower computational cost.
-This indicates that our model offers higher efficiency and better scalability in practical applications.
-The optimization of performance, computational cost, and parameter size makes our model suitable for resource-constrained environments and lays a solid foundation for its future promotion and application in various scenarios.
+To substantiate the approach, we evaluate on the CPCGEA dataset, comparing our model with several strong baselines.
+Figure 2 and Table XIII clearly indicate that MCFNet achieves excellent segmentation performance alongside lower parameter count and computation than competing approaches.
+These results indicate that our model delivers higher computational efficiency and superior scalability in real-world deployments.
+The optimization of performance, computational cost, and parameter size makes our model suitable for resource-constrained environments and lays a solid foundation for its future application in various scenarios.
 V Conclusion
-We design a novel multi-scale, multi-resolution cascade fusion network called MCFNet.
-By capturing rich image features from multi-scale and multi-resolution images, our designed model is better suited to various application scenarios.
-We propose an adaptive loss aggregation strategy called Adaptive-MFA to optimize model training and performance.
-We conduct experiments on multiple datasets covering whole-body risk organs and tumors to fully validate the robustness and generalization capability of the model.
-Experimental results show that our MCFNet exhibits excellent segmentation performance and good generalization ability across multiple datasets.
-In practical clinical applications, the efficient segmentation capability of MCFNet can significantly enhance the accuracy and efficiency of medical image analysis, reduce the workload of doctors, and improve the precision of diagnosis and treatment.
-By accurately segmenting risk organs and tumor regions, MCFNet can assist doctors in better formulating treatment plans, thereby improving patient outcomes.
-Additionally, the good generalization ability of MCFNet means it can adapt to image data generated by different hospitals and equipment, making it a promising tool for widespread use in clinical practice.
-This can drive advancements and progress in the field of medical image analysis.
+We design a novel multi-scale cascaded fusion network called MCFNet.
+This network cascades two backbones, FCB and SEB, to capture rich features from multi-scale and multi-resolution images.
+Among them, we introduce an innovative Linear Attention Transformer module that ensures efficient feature extraction while reducing computational and storage requirements.
+Finally, we propose an adaptive loss aggregation strategy called Adaptive-MFA to optimize model training and performance.
+Extensive experiments on multiple benchmarks demonstrate that MCFNet achieves high segmentation performance with strong generalization.
+MCFNet delivers efficient, high-accuracy segmentation, improving medical image analysis and enhancing clinical precision.
+By accurately segmenting OARs and tumor regions, it assists doctors in formulating better treatment plans, thereby improving patient outcomes.
+Its strong generalization ability allows MCFNet to adapt to data generated by different hospitals and equipment, making it a promising tool for widespread clinical application and advancing medical image analysis.